### v2

# Seeing Through Deception: Uncovering Misleading Creator Intent in Multimodal News with Vision-Language Models

Abstract
The real-world impact of misinformation stems from the underlying misleading narratives that creators seek to convey. As such, interpreting misleading creator intent is essential for multimodal misinformation detection (MMD) systems aimed at effective information governance. In this paper, we introduce an automated framework that simulates real-world multimodal news creation by explicitly modeling creator intent through two components: the desired influence and the execution plan. Using this framework, we construct DeceptionDecoded, a large-scale benchmark comprising 12,000 image-caption pairs aligned with trustworthy reference articles. The dataset captures both misleading and non-misleading intents and spans manipulations across visual and textual modalities. We conduct a comprehensive evaluation of 14 state-of-the-art vision-language models (VLMs) on three intent-centric tasks: (1) misleading intent detection, (2) misleading source attribution, and (3) creator desire inference. Despite recent advances, we observe that current VLMs fall short in recognizing misleading intent, often relying on spurious cues such as superficial cross-modal consistency, stylistic signals, and heuristic authenticity hints. Our findings highlight the pressing need for intent-aware modeling in MMD and open new directions for developing systems capable of deeper reasoning about multimodal misinformation.  [^1] Link to DeceptionDecoded: [GitHub Repo].
Seeing Through Deception: Uncovering Misleading Creator Intent in
Multimodal News with Vision-Language Models
Jiaying Wu1Fanxiao Li2Min-Yen Kan1Bryan Hooi1 1National University of Singapore   2Yunnan University jiayingwu@u.nus.edu, lifanxiao@mail.ynu.edu.cn, {kanmy, bhooi}@comp.nus.edu.sg
Content Warning: this paper contains
potentially harmful text and images.
1 Introduction
Figure 1: The importance of detecting misleading creation intent in multimodal news beyond surface-level cross-modal consistency. For instance, a malicious creator can craft captions that are semantically consistent with the image yet deliberately convey a false narrative.
Multimodal misinformation, which combines persuasive text with compelling visuals, poses significant threats to public understanding and can lead to serious societal harm Zhou and Zafarani (2020); Alam et al. (2022); Do Nascimento et al. (2022). A growing body of research in multimodal misinformation detection (MMD) focuses on identifying cross-modal misalignment, particularly in two major forms: (1) out-of-context (OOC) misinformation Yuan et al. (2023); Qi et al. (2024), where images and captions from unrelated events or time periods are falsely paired; and (2) multimodal media manipulation  Shao et al. (2023); Liu et al. (2025), involving subtle changes such as shifts in facial expressions or caption phrasing that alter the perceived message. However, current MMD benchmarks often simulate such misalignments in artificial ways, either by mismatching image and text with CLIP-based Radford et al. (2021) similarity scores Luo et al. (2021) or through sentiment word substitutions Sudhakar et al. (2019); Shao et al. (2023). These arbitrary strategies create a fundamental gap between constructed benchmarks and the complexity of real-world MMD.
Beyond surface-level cross-modal misalignment, the key to effective real-world misinformation governance lies in detecting and understanding misleading creator intent Appelman et al. (2022); Jaidka et al. (2025). Many misinformation campaigns are deliberately crafted to advance specific agendas, often without the audience’s awareness, yet still manage to substantially influence public opinion Ecker et al. (2022); Broda and Strömbäck (2024). For example, as illustrated in Figure 1, a slight modification to a news caption can lead readers to attribute significant casualties to government negligence, ultimately eroding institutional trust. Although preliminary efforts have explored intent interpretation Da et al. (2021); Gabriel et al. (2022); Wang et al. (2023), they are often limited to unimodal settings and fail to capture the synergy between visual and textual cues. Moreover, these approaches rely on reader inference of creator intent, which are prone to error due to the deceptive framing of such content. These limitations highlight a pressing need for transparent modeling of ground-truth creator intents in multimodal content, along with systematic benchmarks for detecting misleading communicative objectives.
To this end, we establish an intent-guided framework for simulating intent-aware multimodal news reporting. Specifically, we define diverse combinations of desired influence and execution plans, two core components that characterize real-world communication strategies Hallahan et al. (2007); Paul (2011). Building on this, we introduce DeceptionDecoded, a large-scale benchmark designed to reveal misleading creator intents in multimodal news. Constructed upon real-world image-caption pairs from the VisualNews Liu et al. (2021) repository, DeceptionDecoded contains 12,000 instances created with misleading and non-misleading intent, spanning a broad spectrum of misleading intent types and framing strategies. Each instance is paired with a trustworthy reference article to serve as contextual grounding, and the quality of the generated data are further validated through human evaluation (Section 3.4). The benchmark supports evaluation of three intent-centric tasks related to MMD: (1) Misleading Intent Detection, which determines whether a news piece was intentionally crafted to mislead; (2) Misleading Source Attribution, which identifies whether the misleading signal originates from the image or the text; and (3) Creator Desire Inference, which infers the creator’s intended area of societal impact, such as political polarization or public health disruption.
Using DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs). Despite recent advances in multimodal reasoning Hurst et al. (2024); Anthropic (2025c), we find that state-of-the-art VLMs still struggle with intent understanding in the context of multimodal news. For instance, even relatively strong models such as GPT-4o-mini OpenAI (2024) achieve only 18.8% accuracy when misleading intent is subtly expressed through caption manipulation. More concerningly, our analysis reveals that VLMs tend to rely on spurious features when detecting misleading intent, including superficial image–text consistency, stylistic patterns in the text, and heuristic cues suggesting authenticity. This over-reliance leads to significant performance degradation under simple adversarial manipulations, such as stylistic reframing Chen and Shu (2024); Wu et al. (2024) and persuasive prompting attacks Zeng et al. (2024). Our benchmark and empirical analyses establish a foundation for future MMD research that targets deeper semantic reasoning and intent-aware understanding beyond shallow cross-modal consistency patterns.
Benchmark Task Setup Content Modality Creator Intent News Context
Textual Visual
EMU Da et al. (2021) visual manipulation understanding - ✔ (viewer-perceived) -
Fakeddit Nakamura et al. (2020) mixed-source MMD - -
MRF Gabriel et al. (2022) reader perception reasoning - ✔ (reader-perceived) -
NewsINT Wang et al. (2023) news intent interpretation - ✔ (reader-perceived) -
PolitiFact Shu et al. (2020) MMD on social media -
GossipCop Shu et al. (2020) MMD on social media -
NewsCLIPpings Luo et al. (2021) OOC misinformation detection - -
DGM4 Shao et al. (2023) multimedia manipulation detection - -
MMFakebench Liu et al. (2025) mixed-source MMD - -
DeceptionDecoded (Ours) misleading intent detection ✔ (creator-produced)
Table 1: Comparison between DeceptionDecoded and prior benchmarks on misinformation-related tasks.
2 Related Work
2.1 Multimodal Misinformation Detection
With the increasing prevalence of persuasive visual–textual misinformation on online platforms, research has expanded from detecting purely textual misinformation Rashkin et al. (2017); Chen and Shu (2024); Wu et al. (2024) to tackling multimodal misinformation detection (MMD). A growing body of work emphasizes misinformation arising from subtle cross-modal misalignments, such as out-of-context (OOC) misinformation Qi et al. (2024) and multimedia manipulation Shao et al. (2023); Liu et al. (2025). Recent advances leverage VLMs like LLaVA Liu et al. (2023) and GPT-4o Hurst et al. (2024) within retrieval-augmented frameworks, where the model reasons over multimodal inputs using retrieved trustworthy evidence Khaliq et al. (2024); Xuan et al. (2024); Zhou et al. (2024); Braun et al. (2024). While effective in grounding responses, these approaches largely overlook the role of news creation intent, which has been identified as a key driver of misinformation Sharma et al. (2019). Our work builds upon this evidence-based paradigm by evaluating VLMs on their ability to detect misleading communicative intent in multimodal news. We show that even the most advanced models still fall short in this setting, highlighting the need for intent-aware frameworks in MMD.
2.2 Misinformation Benchmarks
Various benchmarks has been developed to support misinformation-related research Hanselowski et al. (2019); Yao et al. (2023). Representative datasets such as PolitiFact and GossipCopShu et al. (2020) collect naturally occurring misinformation on social media. Other works approach MMD from the perspective of multimedia manipulation; for example, NewsCLIPpings Shu et al. (2020) and DGM4 Shao et al. (2023) introduce cross-modal mismatches by substituting images or captions based on visual or textual similarity. Mixed-source datasets such as Fakeddit Nakamura et al. (2020) and MMFakeBench Liu et al. (2025) include both naturally occurring and synthetically manipulated content. While preliminary efforts such as EMU Da et al. (2021), NewsINT Wang et al. (2023), and MRF Gabriel et al. (2022) have considered the role of news creator intent, they are limited to unimodal misinformation. Moreover, the notion of creator intent in these datasets is typically annotated from readers’ subjective perspective, which may misalign with the actual communicative purpose. In this work, we introduce DeceptionDecoded, a benchmark featuring creator-produced multimodal misinformation, where intent is explicitly defined during content generation. A comparison of DeceptionDecoded to prior work is shown in Table 1.
Figure 2: DeceptionDecoded: Overview of multimodal news curation guided by diverse simulated news creator intents, including both misleading and non-misleading cases.
3 DeceptionDecoded
To systematically investigate the role of misleading creator intent in multimodal news, it is crucial to construct a large-scale benchmark that explicitly defines diverse communicative intents across various news domains and uses these intents to guide content generation. However, this is challenging to achieve with real-world news data, where content is already published and annotators can only infer intent retrospectively from a reader’s perspective. To overcome this, we introduce DeceptionDecoded (overviewed in Figure 2), a dataset of 12,000 multimodal news instances generated through an intent-guided synthesis pipeline. Each instance is grounded in a predefined creator intent setup, enabling controlled formulation while preserving content realism. Data quality is validated through rigorous human evaluation.
3.1 Source News Collection
We draw from VisualNews Liu et al. (2021), a large-scale repository of trustworthy multimodal news, to ensure factual grounding. Specifically, we select 10 topic categories that offer broad societal relevance and sample richness. These span diverse domains such as politics, disasters, and public health. The full list of selected topics is provided in Appendix A.1.
To support meaningful evaluation of misinformation detection, we apply rigorous filtering criteria guided by principles from both misinformation governance and professional journalism. From a governance perspective, we prioritize content that serves the public interest Kruger et al. (2024), focusing on events with high societal relevance and potential for wide public impact. In line with journalistic standards, we require that selected content demonstrates professionalism, neutrality, and clarity Maras (2013); Farley et al. (2014), ensuring that language is objective, free from bias, and unambiguous. To further mitigate harm, we exclude instances that reference specific individuals or identifiable public figures (detailed in Appendix A.2).
This process yields 2,000 high-quality samples, evenly distributed across the 10 selected topics. Each instance is structured as $N=\{I,T,A\}$, where $I$ is the news image, $T$ is the accompanying caption, and $A$ is a trusted reference article.
3.2 Creator Intent Establishment
Given the context provided by each original news instance $N=\{I,T,A\}$, we simulate both malicious and trustworthy content creators using GPT-4o Hurst et al. (2024), leveraging its demonstrated capacity for social simulation Anthis et al. (2025).
Inspired by strategic communication theories Hallahan et al. (2007); Paul (2011) and prior work on textual intent inference Wang et al. (2023), we conceptualize creator intent ($C_{\text{int}}$) along two core dimensions: (1) desired influence and (2) execution plan. The desired influence refers to the societal sector(s) the creator aims to affect, selected from a predefined list of options (up to three per instance; see Appendix A.3). The execution plan is generated as open-ended text, detailing how the creator intends to achieve the intended influence.
3.3 Intent-Guided Multimodal News Creation
Guided by $C_{\text{int}}$, we generate misleading and non-misleading variants of each news piece obtained in Section 3.1, by either modifying the image or text. Misleading intents are categorized into two levels: (1) subtle (e.g., minor distortions to framing or background details to nudge reader interpretation) and (2) significant (e.g., major alterations that substantially change the perceived message). For non-misleading cases, we prompt the creator to faithfully paraphrase the caption or reconstruct the image while retaining alignment with the article.
For textual modification, we use GPT-4o Hurst et al. (2024) to generate captions aligned with the specified intent, as defined in Section 3.2, while keeping the original image unchanged. This produces a sample of the form $N_{\text{text}}=\{I,T^{\prime},A\}$, where $I$ is the original image, $T^{\prime}$ is the modified caption, and $A$ is the associated reference article.
For visual modification, we prompt GPT-4o to produce textual descriptions reflecting the intended visual manipulation. These descriptions are then used to generate images via the open-source FLUX.1 [dev] model Black Forest Labs (2024). The resulting sample is of the form $N_{\text{image}}=\{I^{\prime},T,A\}$, where $I^{\prime}$ is the modified image and $T$ is the original caption.
Dataset Statistics.
For each of the 2,000 filtered news samples from Section 3.1, we generate misleading and non-misleading variants for each of the six creator intent categories outlined in Figure 2, resulting in a total of 12,000 instances. All samples are crafted to simulate the professional tone of real-world news reporting. Data examples are provided in Appendix A.4.
3.4 Human Verification of Data Quality
To assess the quality of news samples created in DeceptionDecoded, we conduct a human evaluation by randomly sampling 2% of the dataset, comprising 120 text and 120 image instances covering all six fine-grained intent categories. The goal is to determine whether human annotators perceive the misleading or non-misleading nature of each sample in alignment with the intent simulated by the news creators.
We recruit three graduate students as annotators. Each annotator is presented with 240 label-shuffled pairs of (original, DeceptionDecoded-generated) news pieces, where the generated version may be either misleading (M) or non-misleading (NM). Annotators are instructed to provide a binary label indicating whether the generated version sufficiently deviates from the original such that it could plausibly mislead an average reader. Additional details on the annotation protocol are provided in Appendix B.
We assess data quality using two metrics: (1) accuracy of labels in DeceptionDecoded (misleading/non-misleading) relative to the human-assigned label, based on aggregated human annotations, and (2) inter-annotator agreement, measured by Fleiss’ $\kappa$ Fleiss (1971). The results indicate high reliability and clarity, with 99.2% Acc. and $\kappa=0.877$ for text samples, and 89.2% Acc. with $\kappa=0.703$ for image samples.
3.5 Evaluation Tasks and Metrics
DeceptionDecoded facilitates the evaluation of the following three intent-centric tasks:
  • Task 1: Misleading Intent Detection. Given an input triplet of either $\{I^{\prime},T,A\}$ or $\{I,T^{\prime},A\}$, the goal is to predict whether the instance is misleading (M) or non-misleading (NM).
  • Task 2: Misleading Source Attribution. A 3-way classification task in which the model must attribute the source of misleading intent to either the image or the text modality, or output “NA” if the instance is non-misleading.
  • Task 3: Creator Desire Inference. A multi-label classification task where the model must identify the creator’s intended societal influence, selecting up to three options from a predefined list (see Appendix A.3).
We evaluate Tasks 1 and 2 using classification accuracy, and Task 3 using F1 score.
4 Experiments
Text Image
Model M-Sub M-Sig NM Avg. M-Sub M-Sig NM Avg.
o4-mini I 64.8 (63.9) 91.2 (87.9) 95.4 (95.4) 83.8 (82.4) 25.7 (20.3) 41.8 (33.8) 91.6 (91.6) 53.0 (48.5)
C 72.4 (70.3) 94.0 (92.4) 88.7 (88.7) 85.0 (83.8) 46.3 (33.7) 66.5 (50.3) 78.7 (78.7) 63.8 (54.2)
Claude-3.7-Sonnet I 64.5 (55.0) 90.0 (82.0) 92.8 (92.8) 82.4 (76.6) 46.9 (43.8) 67.2 (64.0) 84.5 (84.5) 66.2 (64.1)
C 67.4 (66.1) 90.2 (89.7) 90.3 (90.3) 82.6 (82.0) 50.3 (31.3) 71.1 (50.6) 81.9 (81.9) 67.8 (54.6)
Gemini-2.5-Pro I 71.0 (70.7) 92.9 (92.5) 93.5 (93.5) 85.8 (85.6) 43.9 (31.1) 66.4 (50.2) 86.4 (86.4) 65.6 (55.9)
C 72.0 (71.7) 93.0 (93.0) 92.2 (92.2) 85.7 (85.6) 53.3 (35.5) 73.9 (51.9) 80.3 (80.3) 69.2 (55.9)
GPT-4o I 67.2 (60.9) 93.1 (85.8) 91.6 (91.6) 84.0 (79.4) 42.6 (39.8) 61.9 (59.9) 85.8 (85.8) 63.4 (61.8)
C 70.2 (63.6) 93.3 (85.1) 86.6 (86.6) 83.4 (78.4) 51.7 (46.9) 69.3 (66.2) 78.0 (78.0) 66.3 (63.7)
Claude-3.5-Sonnet I 58.0 (53.2) 88.5 (84.0) 94.8 (94.8) 80.4 (77.3) 25.3 (18.8) 42.7 (34.0) 93.8 (93.8) 53.9 (48.9)
C 66.8 (66.0) 92.1 (91.6) 90.3 (90.3) 83.1 (82.6) 41.9 (23.7) 63.5 (37.4) 87.0 (87.0) 64.1 (49.4)
Gemini-1.5-Pro I 84.7 (83.6) 97.0 (96.3) 76.5 (76.5) 86.1 (85.5) 54.6 (36.1) 72.4 (51.6) 74.8 (74.8) 67.3 (54.2)
C 80.5 (79.0) 94.6 (94.0) 78.0 (78.0) 84.4 (83.7) 56.9 (34.6) 74.5 (50.1) 72.2 (72.2) 67.9 (52.3)
GPT-4o-mini I 7.3 (5.1) 25.3 (19.8) 99.3 (99.3) 44.0 (41.4) 5.2 (4.6) 9.7 (9.1) 99.2 (99.2) 38.0 (37.6)
C 18.8 (17.2) 45.9 (43.2) 95.7 (95.7) 53.5 (52.0) 21.3 (18.1) 35.8 (33.2) 94.6 (94.6) 50.6 (48.6)
Claude-3.5-Haiku I 3.9 (2.7) 18.2 (12.6) 99.9 (99.9) 40.7 (38.4) 1.1 (0.9) 1.3 (1.2) 99.9 (99.9) 34.1 (34.0)
C 4.2 (3.5) 15.7 (14.1) 99.7 (99.7) 39.9 (39.1) 3.0 (2.7) 5.6 (5.2) 99.7 (99.7) 36.1 (35.9)
Qwen2.5-VL-72B I 47.8 (44.1) 81.0 (75.0) 97.0 (97.0) 75.3 (72.0) 19.5 (17.0) 30.2 (26.9) 92.6 (92.6) 47.4 (45.5)
C 49.6 (47.4) 82.5 (79.9) 94.4 (94.4) 75.5 (73.9) 28.8 (25.3) 42.2 (38.8) 88.6 (88.6) 53.2 (50.9)
Qwen2.5-VL-32B I 18.3 (17.3) 45.4 (44.0) 98.4 (98.4) 54.0 (53.2) 16.0 (15.3) 25.9 (25.2) 96.7 (96.7) 46.2 (45.7)
C 27.5 (27.1) 58.6 (58.1) 97.3 (97.3) 61.1 (60.8) 21.8 (17.9) 33.6 (28.3) 94.6 (94.6) 50.0 (46.9)
Llama-3.2-11B I 8.5 (7.9) 18.1 (13.7) 97.0 (92.7) 41.2 (38.1) 7.0 (4.7) 8.4 (7.1) 96.6 (93.9) 37.3 (35.2)
C 9.1 (5.7) 19.4 (9.2) 94.7 (88.5) 41.1 (34.5) 11.9 (12.2) 15.5 (13.9) 94.2 (88.0) 40.5 (38.0)
InternVL-8B I 3.4 (6.9) 3.9 (6.9) 97.6 (89.4) 35.0 (34.4) 2.3 (7.7) 2.1 (8.8) 98.0 (88.1) 34.1 (34.9)
C 5.5 (2.2) 6.2 (2.4) 96.3 (89.5) 36.0 (31.4) 6.0 (11.4) 7.1 (12.1) 96.3 (89.4) 36.5 (37.6)
LLaVA-v1.6-7B I 0.0 (0.1) 0.0 (0.1) 100.0 (100.0) 33.3 (33.4) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
C 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
Qwen-VL-7B I 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
C 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
Table 2: VLM accuracy (%) on misleading intent detection (black) and misleading source attribution (colored). I refers to the implication-oriented approach, and C refers to the consistency-oriented approach (see Section 4.1). Green cells indicate the top-2 best-performing models. Red cells indicate hallucinations observed in smaller VLMs, where the model attributes a misleading source (M) despite predicting the instance as non-misleading (NM).
Text Image
Model M-Sub M-Sig Avg. M-Sub M-Sig Avg.
o4-mini 60.9 79.7 70.3 37.7 56.3 47.0
Claude-3.7-Sonnet 61.1 82.6 71.9 44.7 65.4 55.1
GPT-4o 57.0 75.9 66.5 39.9 55.4 47.7
Claude-3.5-Sonnet 59.4 82.8 71.1 35.0 55.0 45.0
Gemini-1.5-Pro 68.5 81.9 75.2 46.6 63.6 55.1
GPT-4o-mini 0.8 3.2 2.0 1.7 2.6 2.2
Claude-3.5-Haiku 0.0 0.0 0.0 0.0 0.0 0.0
Qwen2.5-VL-72B 44.1 74.0 59.1 24.6 37.6 31.1
Qwen2.5-VL-32B 0.0 0.0 0.0 0.4 0.0 0.2
Table 3: Performance (F1%) of VLMs on creator desire inference. Green cells indicate the best-performing models, while red cells indicate models that failed to make a single correct prediction.
4.1 Experimental Details
With DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs) spanning a range of sizes, model families, and access levels. These include (1) multimodal Large Reasoning Models (MLRMs), such as Claude-3.7-Sonnet Anthropic (2025c); (2) proprietary multimodal Large Language Models (MLLMs), such as GPT-4o Hurst et al. (2024); and (3) open-source MLLMs, such as the Qwen-2.5-VL series Bai et al. (2025). The complete list of models used in our experiments is provided in Table 7.
We evaluate two representative paradigms in misinformation detection, distinguished by their reasoning focus: (1) The implication-oriented (I) approach, which focuses on the inferred implications of the news sample; and (2) The consistency-oriented (C) approach, which assesses consistency both across modalities and between the image–caption pair and the reference article. Unless otherwise specified, all experiments on DeceptionDecoded default to the consistency-oriented (C) approach. Experimental setup and evaluation prompt details are shown in Appendix C.1.
4.2 Main Results
Consistency-Oriented Reasoning Generally Outperforms Implication-Oriented Reasoning. As shown in Table 2, a possible reason is that misleading news is often constructed through deliberate content manipulation. Detecting unsubstantiated inconsistencies, either between the image and its caption or between the image-caption pair and the reference article, helps models better identify misleading intent.
Model All Misleading (Acc. %) All Non-Misleading (Acc. %)
Original w/ Helpful Hint w/ Adversarial Hint Original w/ Helpful Hint w/ Adversarial Hint
o4-mini 69.8 81.0 (+11.2) 46.8 (-23.0) 83.7 90.2 (+6.5) 68.4 (-15.3)
Claude-3.7-Sonnet 69.7 88.5 (+18.8) 49.3 (-20.4) 86.1 93.9 (+7.8) 58.5 (-27.6)
Gemini-2.5-Pro 73.0 84.6 (+11.6) 43.2 (-29.8) 86.3 99.8 (+13.5) 72.8 (-13.5)
GPT-4o 71.1 87.6 (+16.5) 29.3 (-41.8) 82.3 97.4 (+15.1) 61.1 (-21.2)
Claude-3.5-Sonnet 66.1 74.6 (+8.5) 47.5 (-18.6) 88.6 95.6 (+7.0) 82.4 (-6.2)
Gemini-1.5-Pro 76.6 96.3 (+19.7) 54.3 (-22.3) 75.1 97.3 (+22.2) 51.0 (-24.1)
GPT-4o-mini 30.4 84.9 (+54.5) 6.4 (-24.0) 95.1 99.8 (+4.7) 50.4 (-44.7)
Claude-3.5-Haiku 7.1 34.5 (+27.4) 0.8 (-6.3) 99.7 100.0 (+0.3) 95.7 (-4.0)
Qwen2.5-VL-72B 50.7 81.9 (+31.2) 30.2 (-20.5) 91.5 98.2 (+6.7) 61.6 (-29.9)
Qwen2.5-VL-32B 35.3 99.5 (+64.2) 1.2 (-34.1) 95.9 99.6 (+3.7) 3.4 (-92.5)
Table 4: Misleading intent detection performance of VLMs when exposed to spurious authenticity-related hints, averaged over both misleading and non-misleading cases. Hint construction details are provided in Section 5.1.
Model M-Sub (Text) M-Sig (Text)
I+T T+A Full I+T T+A Full
GPT-4o 21.8 63.4 70.2 51.6 88.4 93.3
GPT-4o-mini 3.6 36.8 18.8 10.2 64.4 45.9
Qwen2.5-72B 10.6 53.6 49.6 24.6 83.8 82.5
Qwen2.5-32B 6.2 35.8 27.5 18.4 67.2 58.6
Table 5: VLM performance (Acc. %) on misleading intent prediction under partial modality settings, suggesting that smaller VLMs may overly rely on image-text consistency for misleading intent detection. (I+T: image & caption only; T+A: text & reference article only; Full: all modalities provided.)
Figure 3: Misleading intent detection performance of VLMs on misleading text presented either in a professional tone (as used in the original DeceptionDecoded setting) or in an explicitly misleading style.
VLMs Struggle to Reason About Misleading Creator Intent.
As shown in Table 2, even state-of-the-art MLRMs such as Claude-3.7 demonstrate limited performance in detecting and attributing misleading creator intent. Performance is even lower on the more complex task of creator desire inference (Table 3), suggesting a deeper challenge in reasoning about communicative goals.
These findings raise alarming concerns. Our setup reflects a realistic deployment scenario, where readers are exposed to subtle yet misleading representations of otherwise trustworthy multimodal news and must rely on automated systems for guidance. The consistent underperformance of VLMs under such conditions suggests an urgent need to investigate the underlying patterns these models exploit when assessing content authenticity, inspiring our in-depth analysis in Section 5.
5 Analysis
5.1 Probing the Limitations of VLMs
VLMs Can Be Misled by Surface-Level Image–Text Consistency.
Although misleading intent is often not substantiated when compared with the trustworthy reference article, the manipulated image and caption typically appear internally consistent, as illustrated in Figure 1. Ideally, models should integrate information across all modalities (i.e., image, text, and article) to detect such deceptive intent. To investigate this, we evaluate VLMs under two partial input settings: (1) Image + Text, and (2) Text + Article.
However, results in Table 5 reveal a contradiction to this ideal. For instance, GPT-4o-mini performs better when detecting misleading text by comparing it with the article, but its performance drops notably when the misleading image is added. This suggests that consistency between image and text can mask deception, misleading models that lack deeper reasoning about cross-source veracity.
VLMs Fail to Detect Misleading Text Framed in Credible-Sounding Styles.
Prior work in textual misinformation detection Chen and Shu (2024); Wu et al. (2024) has shown that large language models (LLMs) often rely on stylistic cues rather than factual correctness when making veracity judgments. To investigate whether this stylistic bias extends to the multimodal setting, we reframe misleading captions in DeceptionDecoded from professional-sounding styles to more overtly deceptive or sensational tones (see Appendix C.2). As illustrated in Figure 3, we observe a similar trend: VLMs are more likely to misclassify misleading content when it is framed in a polished, authoritative style. This suggests that current models may be overly influenced by surface-level linguistic cues, posing a serious risk when misinformation is intentionally crafted to appear credible.
Models React Strongly to Spurious Authenticity Cues in Prompts.
To examine whether VLMs are influenced by high-level framing cues that are not grounded in the input content, we inject two types of authenticity-related hints into the prompt (see Appendix C.3): (1) Skeptical, which presumes the news piece may contain intentional distortions. This hint serves as a helpful signal for misleading samples but becomes adversarial for non-misleading samples; and (2) Trusting, which assumes the content is reliable, reversing the helpful and adversarial roles.
As shown in Table 4, these prompts produce significant and consistent performance changes: they improve accuracy in helpful cases and degrade performance in adversarial ones. This reveals that VLMs are susceptible to spurious prior assumptions about content authenticity, even when such cues are not supported by the input.
5.2 Impact of Image Generation Advances
DeceptionDecoded includes reasonably high-quality images generated using the FLUX model, as supported by human validation results in Section 3.4. As image generation capabilities continue to advance, we conduct a pilot study to assess whether VLM performance improves when exposed to higher-fidelity images generated by GPT-image-1 OpenAI (2025a), using a subset of 200 sampled misleading instances.
As shown in Table 6 and Figure 4, we observe modest performance gains, likely due to the more vivid visual portrayal of creator intent. However, despite these improvements, understanding the underlying communicative intent behind multimodal news remains a significant challenge even when image realism is enhanced. These findings highlight that surface-level fidelity alone is insufficient, reinforcing the need for models that reason explicitly about creator objectives.
The growing realism of deceptive images as reflected by GPT-generated instances introduces a significant risk. When high-fidelity images convincingly align with misleading textual narratives, the deception becomes more subtle and harder to detect. These advances further underscore the importance of reasoning about the creator’s underlying intent, rather than relying solely on content realism or modality alignment.
Figure 4: Case Study: a VLM (GPT-4o-mini) fails to detect misleading creator intent in a FLUX-generated image due to its inability to recognize the unsubstantiated presence of a crowd with fireworks, given the news context. The state-of-the-art GPT-image-1 model produces a clearer depiction, leading to a correct prediction.
Model M-Sub (Image) M-Sig (Image)
Flux GPT-image Flux GPT-image
GPT-4o 51.0 66.0 69.0 84.0
Claude-3.5-Sonnet 42.5 64.0 65.0 80.0
GPT-4o-mini 22.5 47.5 37.5 61.5
Claude-3.5-Haiku 3.0 9.0 6.0 18.0
Qwen2.5-72B 22.5 53.5 33.0 57.0
Qwen2.5-32B 30.0 37.0 46.5 46.0
Table 6: Misleading intent detection performance of VLMs on images created using different image generation models, evaluated on a pilot test set of 200 samples. While performance improves with higher-fidelity GPT images, a substantial gap remains.
6 Conclusion
In this work, we establish the important yet underexplored task of understanding creator intent in multimodal misinformation. We introduce DeceptionDecoded, a large-scale benchmark designed to evaluate VLMs across intent-centric tasks. Through extensive experiments, we uncover alarming vulnerabilities in state-of-the-art VLMs, which over-rely on spurious shortcuts such as cross-modal consistency, stylistic cues, and heuristic indicators of authenticity to detect misleading intent. Our findings pave the way for future research on MMD systems capable of intent-aware, authenticity-centric semantic reasoning toward effective misinformation governance.
Limitations
Through DeceptionDecoded, we establish a framework for understanding misleading creator intent in multimodal news, providing a meaningful first step toward modeling and detecting misleading creator intent. While DeceptionDecoded presents a structured and scalable framework, it also opens up several avenues for further research in advancing intent-aware misinformation governance across more realistic and socially grounded scenarios.
First, although our framework supports the evaluation of creator intent interpretation through tasks such as misleading intent detection and creator desire inference, fully open-ended intent articulation and corresponding evaluation remains an open challenge. Future work could explore adaptive prompting strategies and leverage the emerging LLM-as-a-Judge paradigm Gu et al. (2024) to assess generated intent representations in more flexible and scalable ways.
Second, our current benchmark focuses on intent detection and interpretation, but does not directly measure downstream societal influence. In real-world scenarios, the effectiveness of misinformation depends not only on creator intent but also on how audiences interpret and react to the content. Incorporating social grounding, such as user feedback signals, comment chains, or engagement patterns on social media, would provide richer context for evaluating the influence of misleading content and calibrating model responses accordingly.
Despite these limitations, DeceptionDecoded represents an important contribution toward systematic, intent-aware evaluation of VLMs in real-world misinformation contexts. We hope our benchmark and insights will serve as a solid foundation for future work.
Ethics Statement
The intent-guided simulation of multimodal misinformation in DeceptionDecoded involves strategies that could potentially be misused to generate misleading content using VLMs and open-source image generation models (e.g., FLUX). Nonetheless, given the realistic and challenging nature of detecting misleading creator intent in multimodal news, we believe it is important to transparently acknowledge these risks while reporting empirical findings on the performance and limitations of current state-of-the-art VLMs.
Our benchmark is developed solely to highlight the important yet underexplored aspect of modeling malicious creator intent in MMD, and to advance the understanding of current system limitations.To minimize harm, we explicitly avoid manipulating news content involving specific, identifiable individuals; all simulated content is based on anonymized or synthetic entities. To further mitigate misuse, we will open-source the dataset and evaluation scripts, but refrain from releasing the specific generation prompts that could be repurposed for deceptive use. Data access will be granted only to verified researchers under a binding usage agreement to ensure responsible use. All data collection and usage comply with the terms of service of the underlying models and platforms.
Appendix A DeceptionDecoded Details
A.1 News Topics
VisualNews Liu et al. (2021) provides fine-grained topic annotations for each news article. For DeceptionDecoded, we select 10 major topic categories based on sample abundance, topical diversity, and real-world relevance.
The selected categories are:
‘world’, ‘science_technology’, ‘politics_elections’, ‘environment’, ‘business_economy’, ‘technology’, ‘disaster_accident’, ‘conflict_attack’, ‘us-news’, ‘health_medicine_environment’
These topics form the foundation for generating diverse and realistic multimodal misinformation instances in DeceptionDecoded.
A.2 Source Data Acquisition
In Section 3.1, we obtain high-quality source data for DeceptionDecoded with the prompt outlined in Figure 8.
A.3 Aspects of Desired Influence
We define a set of major desired aspects of societal influence to facilitate creator intent establishment in Section 3.2. The aspects are listed as follows:
Political Polarization, Social Polarization, Cultural and Religious Polarization, Economic Misinformation, Public Health and Safety, Environmental and Scientific Polarization, Geopolitical and International Relations, Psychological and Emotional Manipulation
A.4 DeceptionDecoded Examples
Figure 5: An example in DeceptionDecoded, showcasing a scenario where the creator intentionally misrepresents the news content with a distorted caption to mislead readers.
Figure 6: An example in DeceptionDecoded, showcasing a scenario where the creator intentionally misrepresents the news content with a distorted image to mislead readers.
We present two DeceptionDecoded data examples in Figure 5 (textual modification) and Figure 6 (visual modification).
To prevent misuse, we refrain from disclosing the data generation prompts for Section 3.3 that could be repurposed for deceptive use.
Appendix B Human Evaluation Details
We illustrate the annotation process for human evaluation in Appendix B in Figure 7.
Figure 7: Illustration of human evaluation instructions for validating the quality of DeceptionDecoded.
Appendix C VLM Evaluation Details
C.1 Experimental Setup
With DeceptionDecoded, we systematically evaluate 14 representative VLMs. Detailed model names and versions are provided in Table 7.
For all evaluations, we set the model temperature to 0 to ensure stable outputs, and run each experiment once. Detailed evaluation prompts for the implication- and consistency-based approaches are shown in Appendix C.4.
C.2 Caption Style Reframing Formulation
We adopt the following prompt for the stylistic reframing experiments in Section 5.1 and Table 3.
You are given a factual news caption. Your task is to rewrite it in an unreliable style: that is, make it sound sensational, biased, emotional, or manipulative, while preserving the original content and factual meaning. Do not add or remove any factual information. Only change the tone and style to reflect how misinformation might manipulate readers through language. Original Caption: {caption} Unreliable-style Caption:
C.3 Authenticity-Related Hint Formulation
To investigate how VLMs respond to subtle authenticity cues, we design two types of prompt-level hints aligned with the analysis in Section 5.1. These hints are inserted into the input to simulate prior assumptions about content reliability.
Hint (Skeptical): “Hint: The news piece is likely to contain intentional distortions. Carefully analyze the image and caption to identify any misleading or manipulative elements.”
Hint (Trusting): “Hint: The news piece is sourced from reputable outlets and is therefore unlikely to contain intentional distortions. ”
C.4 Evaluation Prompts
We present our prompts for evaluating VLMs in Figure 9 (implication-oriented) and Figure 10 (consistency-oriented).
Data filtering in DeceptionDecoded You are a senior journalist tasked with evaluating a news caption based on five key criteria: - Public Interest: Does the caption describe a significant event likely to capture public attention? - Professionalism: Does it maintain a professional, objective tone? - Neutrality: Is it free from political, ethnic, gender, or religious bias? - Anonymity: Does it avoid mentioning any specific person’s name or national/world leaders? - Clarity: Is the message clearly conveyed without ambiguity, misunderstanding, or controversy? Conclude with a single word “Yes” if the caption meets all five criteria or a single word “No” if it fails any. Caption: {caption} Response:
Figure 8: Prompt for selecting VisualNews Liu et al. (2021) samples as source data for DeceptionDecoded.
Implication-Oriented Approach (I) Your task is to determine whether a news piece has been intentionally distorted to mislead readers. The news piece consists of an image with its caption: {caption}. You are also provided a trustworthy reference article related to the same event: {article}. Follow the structured steps below and strictly adhere to the output format. Important Note: Focus only on the content of the news piece. The use of AI tools to assist image creation does not, by itself, indicate distortion. Answer the following questions: Q1. Has the news piece been intentionally distorted by its creator to mislead readers, such that it conveys a biased, controversial, or manipulated message? - Respond with one of the following: [Yes, No]. Then provide a brief explanation. Q2. What is the source of the manipulation—text or image? - If answer to Q1 is “Yes”, answer with one of the following: [Text, Image]. - If answer to Q1 is “No”, answer: “NA”. Q3. What does the distortion suggest about the news publisher’s intent? - If answer to Q1 is “Yes” : Primary Influence (Desire): Select up to three from the following list, ranked by relevance: [Political Polarization, Social Polarization, Cultural and Religious Polarization, Economic Misinformation, Public Health and Safety, Environmental and Scientific Polarization, Geopolitical and International Relations, Psychological and Emotional Manipulation]. - If answer to Q1 is “No”, respond with “NA” for Q3. Output Format: Respond in the following exact JSON format, preserving the structure, punctuation, and delimiters. Do not add any extra line breaks or formatting outside this structure: { “pred”: “Answer to Q1, followed by a brief explanation”, “source”: "Answer to Q2", “desire”: "Answer to Q3" }
Figure 9: Prompt for the implication-oriented approach (I) for detecting misleading news creator intent.
Consistency-Oriented Approach (C) Your task is to determine whether a news piece has been intentionally distorted to mislead readers. The news piece consists of an image with its caption: {caption}. You are also provided a trustworthy reference article related to the same event: {article}. Follow the structured steps below and strictly adhere to the output format. Important Note: Focus only on the content of the news piece. The use of AI tools to assist image creation does not, by itself, indicate distortion. Answer the following questions: Q1. Has the news piece been intentionally distorted by its creator to mislead readers, such that there exists inconsistencies between the image and caption, or between the news piece and the reference article? - Respond with one of the following: [Yes, No]. Then provide a brief explanation. Q2. What is the source of the manipulation—text or image? - If answer to Q1 is “Yes”, answer with one of the following: [Text, Image]. - If answer to Q1 is “No”, answer: “NA”. Q3. What does the distortion suggest about the news publisher’s intent? - If answer to Q1 is “Yes” : Primary Influence (Desire): Select up to three from the following list, ranked by relevance: [Political Polarization, Social Polarization, Cultural and Religious Polarization, Economic Misinformation, Public Health and Safety, Environmental and Scientific Polarization, Geopolitical and International Relations, Psychological and Emotional Manipulation]. - If answer to Q1 is “No”, respond with “NA” for Q3. Output Format: Respond in the following exact JSON format, preserving the structure, punctuation, and delimiters. Do not add any extra line breaks or formatting outside this structure: { “pred”: “Answer to Q1, followed by a brief explanation”, “source”: "Answer to Q2", “desire”: "Answer to Q3" }
Figure 10: Prompt for the consistency-oriented approach (C) for detecting misleading news creator intent.

### v3

@@
-Seeing Through Deception: Uncovering Misleading Creator Intent in Multimodal News with Vision-Language Models
+Seeing Through Deception: Uncovering Misleading Creator Intent in Multimodal News with Vision–Language Models
 Abstract
-The real-world impact of misinformation stems from the underlying misleading narratives that creators seek to convey.
-As such, interpreting misleading creator intent is essential for multimodal misinformation detection (MMD) systems aimed at effective information governance.
-In this paper, we introduce an automated framework that simulates real-world multimodal news creation by explicitly modeling creator intent through two components: the desired influence and the execution plan.
-Using this framework, we construct DeceptionDecoded, a large-scale benchmark comprising 12,000 image-caption pairs aligned with trustworthy reference articles.
-The dataset captures both misleading and non-misleading intents and spans manipulations across visual and textual modalities.
-We conduct a comprehensive evaluation of 14 state-of-the-art vision-language models (VLMs) on three intent-centric tasks: (1) misleading intent detection, (2) misleading source attribution, and (3) creator desire inference.
-Despite recent advances, we observe that current VLMs fall short in recognizing misleading intent, often relying on spurious cues such as superficial cross-modal consistency, stylistic signals, and heuristic authenticity hints.
-Our findings highlight the pressing need for intent-aware modeling in MMD and open new directions for developing systems capable of deeper reasoning about multimodal misinformation.
-[^1] Link to DeceptionDecoded: [GitHub Repo].
-Seeing Through Deception: Uncovering Misleading Creator Intent in
-Multimodal News with Vision-Language Models
-Jiaying Wu1Fanxiao Li2Min-Yen Kan1Bryan Hooi1 1National University of Singapore   2Yunnan University jiayingwu@u.nus.edu, lifanxiao@mail.ynu.edu.cn, {kanmy, bhooi}@comp.nus.edu.sg
-Content Warning: this paper contains
-potentially harmful text and images.
+The impact of misinformation arises not only from factual inaccuracies but also from the misleading narratives that creators deliberately embed.
+Interpreting such creator intent is therefore essential for multimodal misinformation detection (MMD) and effective information governance.
+To this end, we introduce DeceptionDecoded, a large-scale benchmark of 12,000 image–caption pairs grounded in trustworthy reference articles, created using an intent-guided simulation framework that models both the desired influence and the execution plan of news creators.
+The dataset captures both misleading and non-misleading cases, spanning manipulations across visual and textual modalities, and supports three intent-centric tasks: (1) misleading intent detection, (2) misleading source attribution, and (3) creator desire inference.
+We evaluate 14 state-of-the-art vision–language models (VLMs) and find that they struggle with intent reasoning, often relying on shallow cues such as surface-level alignment, stylistic polish, or heuristic authenticity signals.
+These results highlight the limitations of current VLMs and position DeceptionDecoded as a foundation for developing intent-aware models that go beyond shallow cues in MMD.
+Content Warning: this paper contains potentially harmful text and images.
 1 Introduction
-Figure 1: The importance of detecting misleading creation intent in multimodal news beyond surface-level cross-modal consistency.
-For instance, a malicious creator can craft captions that are semantically consistent with the image yet deliberately convey a false narrative.
-Multimodal misinformation, which combines persuasive text with compelling visuals, poses significant threats to public understanding and can lead to serious societal harm Zhou and Zafarani (2020); Alam et al.
-(2022); Do Nascimento et al.
-(2022).
-A growing body of research in multimodal misinformation detection (MMD) focuses on identifying cross-modal misalignment, particularly in two major forms: (1) out-of-context (OOC) misinformation Yuan et al.
-(2023); Qi et al.
-(2024), where images and captions from unrelated events or time periods are falsely paired; and (2) multimodal media manipulation  Shao et al.
-(2023); Liu et al.
-(2025), involving subtle changes such as shifts in facial expressions or caption phrasing that alter the perceived message.
-However, current MMD benchmarks often simulate such misalignments in artificial ways, either by mismatching image and text with CLIP-based Radford et al.
-(2021) similarity scores Luo et al.
-(2021) or through sentiment word substitutions Sudhakar et al.
-(2019); Shao et al.
-(2023).
-These arbitrary strategies create a fundamental gap between constructed benchmarks and the complexity of real-world MMD.
-Beyond surface-level cross-modal misalignment, the key to effective real-world misinformation governance lies in detecting and understanding misleading creator intent Appelman et al.
-(2022); Jaidka et al.
-(2025).
-Many misinformation campaigns are deliberately crafted to advance specific agendas, often without the audience’s awareness, yet still manage to substantially influence public opinion Ecker et al.
-(2022); Broda and Strömbäck (2024).
-For example, as illustrated in Figure 1, a slight modification to a news caption can lead readers to attribute significant casualties to government negligence, ultimately eroding institutional trust.
-Although preliminary efforts have explored intent interpretation Da et al.
-(2021); Gabriel et al.
-(2022); Wang et al.
-(2023), they are often limited to unimodal settings and fail to capture the synergy between visual and textual cues.
+Multimodal misinformation, which combines persuasive text with compelling visuals, poses significant threats to public understanding and can lead to serious societal harm (Zhou & Zafarani, 2020; Alam et al., 2022; Do Nascimento et al., 2022).
+Research in multimodal misinformation detection (MMD) has primarily focused on cross-modal misalignment, typically in two forms: (1) out-of-context (OOC) misinformation (Yuan et al., 2023; Qi et al., 2024), where images and captions from unrelated events or time periods are falsely paired; and (2) multimodal media manipulation (Shao et al., 2023; Liu et al., 2025), involving subtle changes to visuals or captions that alter interpretation.
+Yet, current benchmarks rely on heuristic strategies such as CLIP-based mismatching (Luo et al., 2021) or sentiment substitutions (Sudhakar et al., 2019; Shao et al., 2023), which fail to capture the complexity of real-world multimodal misinformation.
+Beyond surface-level cross-modal misalignment, the key to effective real-world misinformation governance lies in detecting and understanding misleading creator intent (Appelman et al., 2022; Jaidka et al., 2025).
+Many misinformation campaigns are deliberately crafted to advance specific agendas, often without the audience’s awareness, yet still manage to substantially influence public opinion (Ecker et al., 2022; Broda & Strömbäck, 2024).
+As illustrated in Figure 1, even slight caption manipulations can significantly distort a trustworthy news context (e.g., by misattributing iceberg collapse to covert military actions) which ultimately erodes institutional trust.
+Although preliminary efforts have explored intent interpretation (Da et al., 2021; Gabriel et al., 2022; Wang et al., 2025), they are often limited to unimodal settings and fail to capture the synergy between visual and textual cues.
 Moreover, these approaches rely on reader inference of creator intent, which are prone to error due to the deceptive framing of such content.
-These limitations highlight a pressing need for transparent modeling of ground-truth creator intents in multimodal content, along with systematic benchmarks for detecting misleading communicative objectives.
-To this end, we establish an intent-guided framework for simulating intent-aware multimodal news reporting.
-Specifically, we define diverse combinations of desired influence and execution plans, two core components that characterize real-world communication strategies Hallahan et al.
-(2007); Paul (2011).
-Building on this, we introduce DeceptionDecoded, a large-scale benchmark designed to reveal misleading creator intents in multimodal news.
-Constructed upon real-world image-caption pairs from the VisualNews Liu et al.
-(2021) repository, DeceptionDecoded contains 12,000 instances created with misleading and non-misleading intent, spanning a broad spectrum of misleading intent types and framing strategies.
-Each instance is paired with a trustworthy reference article to serve as contextual grounding, and the quality of the generated data are further validated through human evaluation (Section 3.4).
-The benchmark supports evaluation of three intent-centric tasks related to MMD: (1) Misleading Intent Detection, which determines whether a news piece was intentionally crafted to mislead; (2) Misleading Source Attribution, which identifies whether the misleading signal originates from the image or the text; and (3) Creator Desire Inference, which infers the creator’s intended area of societal impact, such as political polarization or public health disruption.
-Using DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs).
-Despite recent advances in multimodal reasoning Hurst et al.
-(2024); Anthropic (2025c), we find that state-of-the-art VLMs still struggle with intent understanding in the context of multimodal news.
-For instance, even relatively strong models such as GPT-4o-mini OpenAI (2024) achieve only 18.8% accuracy when misleading intent is subtly expressed through caption manipulation.
-More concerningly, our analysis reveals that VLMs tend to rely on spurious features when detecting misleading intent, including superficial image–text consistency, stylistic patterns in the text, and heuristic cues suggesting authenticity.
-This over-reliance leads to significant performance degradation under simple adversarial manipulations, such as stylistic reframing Chen and Shu (2024); Wu et al.
-(2024) and persuasive prompting attacks Zeng et al.
-(2024).
-Our benchmark and empirical analyses establish a foundation for future MMD research that targets deeper semantic reasoning and intent-aware understanding beyond shallow cross-modal consistency patterns.
-Benchmark Task Setup Content Modality Creator Intent News Context
+These limitations highlight a pressing need for transparent modeling of ground-truth creator intents in multimodal content, along with systematic evaluation frameworks for detecting misleading communicative objectives.
+Figure 1: Why creator intent detection matters in multimodal news.
+Creators can distort trustworthy news contexts by crafting intent-loaded misinformation that appears semantically aligned (e.g., portraying broken icebergs) yet deliberately conveys false narratives (e.g., attributing them to secret underwater nuclear tests).
+To address this gap, we introduce an intent-guided framework for simulating intent-aware multimodal news reporting.
+In this framework, intent is defined as a combination of desired influence and execution plan, two components grounded in communication strategy theory (Hallahan et al., 2007; Paul, 2011).
+Leveraging this formulation, we construct DeceptionDecoded, a large-scale benchmark of 12,000 image–caption–article triplets $(I,T,A)$, where $I$ is the news image, $T$ is the accompanying caption, and $A$ is a trustworthy reference article (i.e., “trustworthy news context”) drawn from the VisualNews dataset of verified reporting (Liu et al., 2021).
+By anchoring manipulations to verifiable news contexts, this design enables explicit labeling of misleading versus non-misleading reporting, while also ensuring coverage of ten high-impact domains selected using journalistic and public-interest criteria.
+The benchmark further models realistic malicious influence scenarios such as political polarization and public health misinformation (Section 3.1), all while preserving a professional news reporting tone.
+Data quality and realism are validated through systematic human evaluation (Section 3.4), and the benchmark supports three intent-centric tasks central to effective MMD: (1) Misleading Intent Detection, which determines whether a multimodal news piece was deliberately crafted to mislead readers; (2) Misleading Source Attribution, which identifies whether the misleading signal originates from the image or the text; and (3) Creator Desire Inference, which infers the targeted area of societal impact (e.g., political polarization or public health disruption).
+Using DeceptionDecoded, we benchmark 14 representative VLMs.
+Despite recent progress in multimodal reasoning (Hurst et al., 2024; Anthropic, 2025b), we find that state-of-the-art VLMs still struggle to recognize misleading creator intent in multimodal news.
+For instance, GPT-4o-mini achieves only 18.8% accuracy when intent is subtly encoded through caption manipulation (OpenAI, 2024).
+Models often rely on superficial cues such as image–text alignment, stylistic polish, or heuristic authenticity signals, leaving them vulnerable to simple adversarial manipulations including stylistic reframing (Chen & Shu, 2024; Wu et al., 2024) and persuasive prompting attacks (Zeng et al., 2024).
+Our analysis further reveals two broader insights.
+First, fine-tuning on DeceptionDecoded transfers improvements to external MMD benchmarks, demonstrating general utility.
+Second, advances in image generation improve visual fidelity, making manipulations more vivid, but VLMs still struggle with intent reasoning even under clearer visuals.
+At the same time, the scalability and realism of ever-evolving generators heighten the risk of persuasive misinformation, positioning DeceptionDecoded as both a principled benchmark and a stress test for developing intent-aware detectors that move beyond surface-level consistency.
+Benchmark Task Setup Content Modality Creator Intent Trustworthy Context
 Textual Visual
@@
 2 Related Work
-2.1 Multimodal Misinformation Detection
-With the increasing prevalence of persuasive visual–textual misinformation on online platforms, research has expanded from detecting purely textual misinformation Rashkin et al.
-(2017); Chen and Shu (2024); Wu et al.
-(2024) to tackling multimodal misinformation detection (MMD).
-A growing body of work emphasizes misinformation arising from subtle cross-modal misalignments, such as out-of-context (OOC) misinformation Qi et al.
-(2024) and multimedia manipulation Shao et al.
-(2023); Liu et al.
-(2025).
-Recent advances leverage VLMs like LLaVA Liu et al.
-(2023) and GPT-4o Hurst et al.
-(2024) within retrieval-augmented frameworks, where the model reasons over multimodal inputs using retrieved trustworthy evidence Khaliq et al.
-(2024); Xuan et al.
-(2024); Zhou et al.
-(2024); Braun et al.
-(2024).
-While effective in grounding responses, these approaches largely overlook the role of news creation intent, which has been identified as a key driver of misinformation Sharma et al.
-(2019).
+2.1 Multimodal Misinformation Detection (MMD)
+With the increasing prevalence of persuasive visual–textual misinformation on online platforms, research has expanded from detecting purely textual misinformation (Rashkin et al., 2017; Chen & Shu, 2024; Wu et al., 2024) to tackling multimodal misinformation detection (MMD).
+A growing body of work emphasizes misinformation arising from subtle cross-modal misalignments, such as out-of-context (OOC) misinformation (Qi et al., 2024) and multimedia manipulation (Shao et al., 2023; Liu et al., 2025).
+Recent advances leverage VLMs like LLaVA (Liu et al., 2023) and GPT-4o (Hurst et al., 2024) within retrieval-augmented frameworks, where the model reasons over multimodal inputs using retrieved trustworthy evidence (Khaliq et al., 2024; Xuan et al., 2024; Zhou et al., 2024; Braun et al., 2024).
+While effective in grounding responses, these approaches largely overlook the role of news creation intent, which has been identified as a key driver of misinformation (Sharma et al., 2019).
 Our work builds upon this evidence-based paradigm by evaluating VLMs on their ability to detect misleading communicative intent in multimodal news.
@@
 2.2 Misinformation Benchmarks
-Various benchmarks has been developed to support misinformation-related research Hanselowski et al.
-(2019); Yao et al.
-(2023).
-Representative datasets such as PolitiFact and GossipCopShu et al.
-(2020) collect naturally occurring misinformation on social media.
-Other works approach MMD from the perspective of multimedia manipulation; for example, NewsCLIPpings Shu et al.
-(2020) and DGM4 Shao et al.
-(2023) introduce cross-modal mismatches by substituting images or captions based on visual or textual similarity.
-Mixed-source datasets such as Fakeddit Nakamura et al.
-(2020) and MMFakeBench Liu et al.
-(2025) include both naturally occurring and synthetically manipulated content.
-While preliminary efforts such as EMU Da et al.
-(2021), NewsINT Wang et al.
-(2023), and MRF Gabriel et al.
-(2022) have considered the role of news creator intent, they are limited to unimodal misinformation.
-Moreover, the notion of creator intent in these datasets is typically annotated from readers’ subjective perspective, which may misalign with the actual communicative purpose.
-In this work, we introduce DeceptionDecoded, a benchmark featuring creator-produced multimodal misinformation, where intent is explicitly defined during content generation.
-A comparison of DeceptionDecoded to prior work is shown in Table 1.
-Figure 2: DeceptionDecoded: Overview of multimodal news curation guided by diverse simulated news creator intents, including both misleading and non-misleading cases.
+A variety of benchmarks have been developed to support misinformation-related research (Hanselowski et al., 2019; Yao et al., 2023).
+Representative datasets such as PolitiFact and GossipCop (Shu et al., 2020) collect naturally occurring misinformation from social media, while others approach multimodal misinformation detection (MMD) through multimedia manipulation; for example, NewsCLIPpings (Luo et al., 2021) and DGM4 (Shao et al., 2023) introduce cross-modal mismatches by substituting images or captions based on visual or textual similarity.
+Mixed-source datasets such as Fakeddit (Nakamura et al., 2020) and MMFakeBench (Liu et al., 2025) combine naturally occurring content with synthetically manipulated samples.
+Preliminary efforts such as EMU (Da et al., 2021), NewsINT (Wang et al., 2025), and MRF (Gabriel et al., 2022) have begun to consider news creator intent, but these remain limited to unimodal settings and typically define intent from readers’ subjective perspectives, which may diverge from the actual communicative purpose of the creator.
+In contrast, we introduce DeceptionDecoded, a benchmark of creator-produced multimodal misinformation where intent is explicitly defined during content generation.
+A comparison of DeceptionDecoded with prior datasets is shown in Table 1.
+Beyond this, DeceptionDecoded is guided by established communication theories, grounded in trustworthy news contexts, and validated through human evaluation, providing both a principled framework for modeling creator intent and a resource with broader utility for advancing general intent-aware MMD.
+Figure 2: DeceptionDecoded: Overview of multimodal news curation guided by diverse simulated creator intents, covering both misleading and non-misleading cases.
 3 DeceptionDecoded
-To systematically investigate the role of misleading creator intent in multimodal news, it is crucial to construct a large-scale benchmark that explicitly defines diverse communicative intents across various news domains and uses these intents to guide content generation.
-However, this is challenging to achieve with real-world news data, where content is already published and annotators can only infer intent retrospectively from a reader’s perspective.
-To overcome this, we introduce DeceptionDecoded (overviewed in Figure 2), a dataset of 12,000 multimodal news instances generated through an intent-guided synthesis pipeline.
-Each instance is grounded in a predefined creator intent setup, enabling controlled formulation while preserving content realism.
-Data quality is validated through rigorous human evaluation.
+To systematically investigate the role of misleading creator intent in multimodal news, it is necessary to construct a large-scale benchmark that (1) explicitly defines diverse communicative intents across multiple domains and (2) uses these intents to guide multimodal content generation.
+Achieving this with real-world news data is difficult, since published articles already reflect finalized narratives and annotators can only infer intent retrospectively from a reader’s perspective.
+To address this challenge, we introduce DeceptionDecoded (Figure 2), a benchmark of 12,000 multimodal news instances created through an intent-guided synthesis pipeline.
+Each instance is anchored in a trustworthy news context drawn from verified articles and paired with a predefined creator intent configuration, which enables controlled manipulation while maintaining content realism.
+The quality and plausibility of the generated data are further validated through rigorous human evaluation (see Section 3.4).
 3.1 Source News Collection
-We draw from VisualNews Liu et al.
-(2021), a large-scale repository of trustworthy multimodal news, to ensure factual grounding.
-Specifically, we select 10 topic categories that offer broad societal relevance and sample richness.
-These span diverse domains such as politics, disasters, and public health.
-The full list of selected topics is provided in Appendix A.1.
-To support meaningful evaluation of misinformation detection, we apply rigorous filtering criteria guided by principles from both misinformation governance and professional journalism.
-From a governance perspective, we prioritize content that serves the public interest Kruger et al.
-(2024), focusing on events with high societal relevance and potential for wide public impact.
-In line with journalistic standards, we require that selected content demonstrates professionalism, neutrality, and clarity Maras (2013); Farley et al.
-(2014), ensuring that language is objective, free from bias, and unambiguous.
-To further mitigate harm, we exclude instances that reference specific individuals or identifiable public figures (detailed in Appendix A.2).
-This process yields 2,000 high-quality samples, evenly distributed across the 10 selected topics.
-Each instance is structured as $N=\{I,T,A\}$, where $I$ is the news image, $T$ is the accompanying caption, and $A$ is a trusted reference article.
+We draw source data from VisualNews (Liu et al., 2021), a large-scale repository of trustworthy multimodal news, to ensure factual grounding.
+From this corpus, we select ten topic categories that provide broad societal relevance and sufficient sample richness, spanning domains such as politics, disasters, and public health.
+The full list of topics is provided in Appendix D.1.
+To support meaningful evaluation of misinformation detection, we apply rigorous filtering criteria informed by both misinformation governance and professional journalism.
+From a governance perspective, we prioritize content that serves the public interest (Kruger et al., 2024), focusing on events with high societal relevance and potential for broad impact.
+Consistent with journalistic standards, we require that selected content demonstrates professionalism, neutrality, and clarity (Maras, 2013; Farley et al., 2014), ensuring that language remains objective, unbiased, and unambiguous.
+To further mitigate harm, we exclude instances referencing specific individuals or identifiable public figures (see Appendix D.2).
+This process yields 2,000 high-quality news samples, evenly distributed across the ten categories.
+Each instance is represented as $N=\{I,T,A\}$, where $I$ is the news image, $T$ is the accompanying caption, and $A$ is the corresponding trusted reference article (i.e., “trustworthy news context”).
 3.2 Creator Intent Establishment
-Given the context provided by each original news instance $N=\{I,T,A\}$, we simulate both malicious and trustworthy content creators using GPT-4o Hurst et al.
-(2024), leveraging its demonstrated capacity for social simulation Anthis et al.
-(2025).
-Inspired by strategic communication theories Hallahan et al.
-(2007); Paul (2011) and prior work on textual intent inference Wang et al.
-(2023), we conceptualize creator intent ($C_{\text{int}}$) along two core dimensions: (1) desired influence and (2) execution plan.
-The desired influence refers to the societal sector(s) the creator aims to affect, selected from a predefined list of options (up to three per instance; see Appendix A.3).
-The execution plan is generated as open-ended text, detailing how the creator intends to achieve the intended influence.
+Given the context provided by each original news instance $N=\{I,T,A\}$, we simulate both malicious and trustworthy content creators using GPT-4o (Hurst et al., 2024), leveraging its demonstrated capacity for social simulation (Anthis et al., 2025).
+Inspired by strategic communication theories (Hallahan et al., 2007; Paul, 2011) and prior work on textual intent inference (Wang et al., 2025), we conceptualize creator intent ($C_{\text{int}}$) along two core dimensions: (1) desired influence and (2) execution plan.
+The desired influence specifies the societal sector(s) the creator seeks to affect (e.g., “public health and safety”), selected from a predefined taxonomy (up to three per instance; see Appendix D.3).
+The execution plan is expressed as open-ended text describing how the creator aims to achieve the intended influence.
 3.3 Intent-Guided Multimodal News Creation
-Guided by $C_{\text{int}}$, we generate misleading and non-misleading variants of each news piece obtained in Section 3.1, by either modifying the image or text.
-Misleading intents are categorized into two levels: (1) subtle (e.g., minor distortions to framing or background details to nudge reader interpretation) and (2) significant (e.g., major alterations that substantially change the perceived message).
-For non-misleading cases, we prompt the creator to faithfully paraphrase the caption or reconstruct the image while retaining alignment with the article.
-For textual modification, we use GPT-4o Hurst et al.
-(2024) to generate captions aligned with the specified intent, as defined in Section 3.2, while keeping the original image unchanged.
-This produces a sample of the form $N_{\text{text}}=\{I,T^{\prime},A\}$, where $I$ is the original image, $T^{\prime}$ is the modified caption, and $A$ is the associated reference article.
-For visual modification, we prompt GPT-4o to produce textual descriptions reflecting the intended visual manipulation.
-These descriptions are then used to generate images via the open-source FLUX.1 [dev] model Black Forest Labs (2024).
-The resulting sample is of the form $N_{\text{image}}=\{I^{\prime},T,A\}$, where $I^{\prime}$ is the modified image and $T$ is the original caption.
+Guided by $C_{\text{int}}$, we generate both misleading and non-misleading variants of each news piece obtained in Section 3.1 by modifying either the image or the text.
+Misleading variants are categorized into two levels: (1) subtle, involving minor distortions to framing or background details that nudge reader interpretation, and (2) significant, involving major alterations that substantially change the perceived message.
+For non-misleading cases, the creator is prompted to faithfully paraphrase the caption or reconstruct the image while retaining consistency with the trustworthy news context $A$.
+For textual modification, we use GPT-4o (Hurst et al., 2024) to generate captions aligned with the specified intent, as defined in Section 3.2, while keeping the original image unchanged.
+This produces samples of the form $N_{\text{text}}=\{I,T^{\prime},A\}$, where $I$ is the original image, $T^{\prime}$ is the modified caption, and $A$ is the corresponding trustworthy news context.
+For visual modification, GPT-4o is prompted to generate textual descriptions of the intended manipulation, which are then used to synthesize images with the open-source FLUX.1 [dev] model (Black Forest Labs, 2024).
+The resulting sample takes the form $N_{\text{image}}=\{I^{\prime},T,A\}$, where $I^{\prime}$ is the modified image, $T$ is the original caption, and $A$ is the trustworthy news context.
 Dataset Statistics.
-For each of the 2,000 filtered news samples from Section 3.1, we generate misleading and non-misleading variants for each of the six creator intent categories outlined in Figure 2, resulting in a total of 12,000 instances.
-All samples are crafted to simulate the professional tone of real-world news reporting.
-Data examples are provided in Appendix A.4.
+For each of the 2,000 filtered news samples from Section 3.1, we generate both misleading and non-misleading variants for each of the six creator intent categories outlined in Figure 2, yielding a total of 12,000 instances.
+All samples are crafted to reflect the professional tone of real-world news reporting.
+Illustrative data examples are provided in Appendix D.4.
 3.4 Human Verification of Data Quality
-To assess the quality of news samples created in DeceptionDecoded, we conduct a human evaluation by randomly sampling 2% of the dataset, comprising 120 text and 120 image instances covering all six fine-grained intent categories.
-The goal is to determine whether human annotators perceive the misleading or non-misleading nature of each sample in alignment with the intent simulated by the news creators.
+To assess the quality of news samples created in DeceptionDecoded through intent-guided creation, we conduct a human evaluation by randomly sampling 2% of the dataset, comprising 120 text-modified and 120 image-modified instances that span all six fine-grained intent categories.
+The primary goal is to verify whether human annotators perceive the misleading or non-misleading nature of each sample in alignment with the intent simulated by the news creators.
 We recruit three graduate students as annotators.
 Each annotator is presented with 240 label-shuffled pairs of (original, DeceptionDecoded-generated) news pieces, where the generated version may be either misleading (M) or non-misleading (NM).
-Annotators are instructed to provide a binary label indicating whether the generated version sufficiently deviates from the original such that it could plausibly mislead an average reader.
-Additional details on the annotation protocol are provided in Appendix B.
-We assess data quality using two metrics: (1) accuracy of labels in DeceptionDecoded (misleading/non-misleading) relative to the human-assigned label, based on aggregated human annotations, and (2) inter-annotator agreement, measured by Fleiss’ $\kappa$ Fleiss (1971).
-The results indicate high reliability and clarity, with 99.2% Acc.
-and $\kappa=0.877$ for text samples, and 89.2% Acc.
-with $\kappa=0.703$ for image samples.
+Annotators are instructed to provide a binary label indicating whether the generated version deviates from the original in a way that could plausibly mislead an average reader.
+Details of the annotation protocol are provided in Appendix E.
+We evaluate data quality using two metrics: (1) accuracy, computed as agreement between the dataset labels and aggregated human annotations, and (2) inter-annotator agreement, measured by Fleiss’ $\kappa$ (Fleiss, 1971).
+Results show high reliability: 99.2% accuracy with $\kappa=0.877$ for text samples and 89.2% accuracy with $\kappa=0.703$ for image samples.
+Beyond verifying the correctness of these binary labels, we also assess data realism and intent alignment.
+As described in Appendix E.2, annotators judged whether each synthesized piece could plausibly appear in today’s news or social media.
+Accuracy exceeded 98% for text and 91% for images in standalone plausibility (judging each piece in isolation), and remained above 87% in pairwise plausibility (judging original $\rightarrow$ manipulated pairs), confirming the contextual realism of the data.
+Furthermore, Appendix E.3 reports evaluations of whether generated manipulations remain faithful to their simulated desired influence and execution plan.
+Accuracy exceeded 92% for text and 86% for images across both dimensions, with substantial inter-annotator agreement.
+Together, these evaluations demonstrate that DeceptionDecoded produces synthetic news samples that are not only correctly labeled as misleading or non-misleading, but also realistic in presentation and meaningfully aligned with the intended creator objectives.
 3.5 Evaluation Tasks and Metrics
-DeceptionDecoded facilitates the evaluation of the following three intent-centric tasks:
+We design three intent-centric tasks in DeceptionDecoded to evaluate complementary dimensions of creator intent reasoning.
   • Task 1: Misleading Intent Detection.
-Given an input triplet of either $\{I^{\prime},T,A\}$ or $\{I,T^{\prime},A\}$, the goal is to predict whether the instance is misleading (M) or non-misleading (NM).
+Given an input triplet of either $\{I^{\prime},T,A\}$ or $\{I,T^{\prime},A\}$, the objective is to predict whether the instance is misleading (M) or non-misleading (NM).
   • Task 2: Misleading Source Attribution.
-A 3-way classification task in which the model must attribute the source of misleading intent to either the image or the text modality, or output “NA” if the instance is non-misleading.
+A three-way classification task in which the model must attribute the source of misleading intent to either the image or the text modality, or output “NA” if the instance is non-misleading.
   • Task 3: Creator Desire Inference.
-A multi-label classification task where the model must identify the creator’s intended societal influence, selecting up to three options from a predefined list (see Appendix A.3).
-We evaluate Tasks 1 and 2 using classification accuracy, and Task 3 using F1 score.
+A multi-label classification task where the model identifies the creator’s intended societal influence, selecting up to three options from a predefined list (see details in Appendix D.3).
+Evaluation for Tasks 1 and 2 is based on classification accuracy, and for Task 3 on F1 score.
 4 Experiments
@@
 C 27.5 (27.1) 58.6 (58.1) 97.3 (97.3) 61.1 (60.8) 21.8 (17.9) 33.6 (28.3) 94.6 (94.6) 50.0 (46.9)
-Llama-3.2-11B I 8.5 (7.9) 18.1 (13.7) 97.0 (92.7) 41.2 (38.1) 7.0 (4.7) 8.4 (7.1) 96.6 (93.9) 37.3 (35.2)
+Llama-3.2-Vision-11B I 8.5 (7.9) 18.1 (13.7) 97.0 (92.7) 41.2 (38.1) 7.0 (4.7) 8.4 (7.1) 96.6 (93.9) 37.3 (35.2)
 C 9.1 (5.7) 19.4 (9.2) 94.7 (88.5) 41.1 (34.5) 11.9 (12.2) 15.5 (13.9) 94.2 (88.0) 40.5 (38.0)
@@
 C 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
-Qwen-VL-7B I 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
+Qwen2.5-VL-7B I 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
 C 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3) 0.0 (0.0) 0.0 (0.0) 100.0 (100.0) 33.3 (33.3)
-Table 2: VLM accuracy (%) on misleading intent detection (black) and misleading source attribution (colored).
-I refers to the implication-oriented approach, and C refers to the consistency-oriented approach (see Section 4.1).
-Green cells indicate the top-2 best-performing models.
-Red cells indicate hallucinations observed in smaller VLMs, where the model attributes a misleading source (M) despite predicting the instance as non-misleading (NM).
+Table 2: VLM accuracy (%) on misleading intent detection (black) and source attribution (colored).
+I denotes the implication-oriented approach, and C the consistency-oriented approach (see Section 4.1).
+Green cells mark the top-2 models.
+Red cells highlight hallucinations in smaller VLMs, where a misleading source (M) is attributed despite the instance being predicted as non-misleading (NM).
+4.1 Experimental Setup
+Using DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs) spanning a range of sizes, model families, and access levels.
+These include (1) multimodal Large Reasoning Models (MLRMs), such as Claude-3.7-Sonnet (Anthropic, 2025b); (2) proprietary multimodal Large Language Models (MLLMs), such as GPT-4o (Hurst et al., 2024); and (3) open-source MLLMs, such as the Qwen-2.5-VL series (Bai et al., 2025).
+The complete list of models used in our experiments is provided in Table 8.
+We evaluate two paradigms in misinformation detection, distinguished by their reasoning focus: (1) The implication-oriented (I) approach, which focuses on the inferred implications of the news sample; and (2) The consistency-oriented (C) approach, which assesses consistency both across modalities and between the image–caption pair and the trustworthy news context.
+Unless otherwise specified, all experiments on DeceptionDecoded default to the consistency-oriented (C) approach.
+Details of the experimental setup and evaluation prompts are provided in Appendix F.1.
+4.2 Main Results
+Consistency-Oriented Reasoning Outperforms Implication-Oriented Reasoning.
+As shown in Table 2, the consistency-oriented approach generally yields stronger performance than the implication-oriented approach.
+A likely explanation is that misleading news is often created through deliberate content manipulation.
+Detecting unsubstantiated inconsistencies—either between the image and its caption or between the image–caption pair and the trustworthy news context—provides models with a more reliable signal for identifying misleading intent.
+VLMs Struggle to Reason About Misleading Creator Intent.
+As shown in Table 2, even state-of-the-art MLRMs such as Claude-3.7 exhibit limited performance in detecting and attributing misleading creator intent.
+Performance drops further on the more challenging task of creator desire inference (Table 4), indicating a deeper difficulty in reasoning about underlying communicative goals.
+These findings raise concerning implications.
+Our evaluation reflects realistic deployment scenarios in which readers encounter subtle yet misleading representations of otherwise trustworthy multimodal news and must rely on automated systems for support.
+The consistent underperformance of VLMs under such conditions highlights the urgent need to examine the features and heuristics these models rely on when assessing authenticity, motivating our in-depth analyses presented in Section 4.3.
 Text Image
@@
 GPT-4o 57.0 75.9 66.5 39.9 55.4 47.7
-Claude-3.5-Sonnet 59.4 82.8 71.1 35.0 55.0 45.0
 Gemini-1.5-Pro 68.5 81.9 75.2 46.6 63.6 55.1
@@
 Qwen2.5-VL-32B 0.0 0.0 0.0 0.4 0.0 0.2
-Table 3: Performance (F1%) of VLMs on creator desire inference. Green cells indicate the best-performing models, while red cells indicate models that failed to make a single correct prediction.
-4.1 Experimental Details
-With DeceptionDecoded, we evaluate 14 representative vision-language models (VLMs) spanning a range of sizes, model families, and access levels.
-These include (1) multimodal Large Reasoning Models (MLRMs), such as Claude-3.7-Sonnet Anthropic (2025c); (2) proprietary multimodal Large Language Models (MLLMs), such as GPT-4o Hurst et al.
-(2024); and (3) open-source MLLMs, such as the Qwen-2.5-VL series Bai et al.
-(2025).
-The complete list of models used in our experiments is provided in Table 7.
-We evaluate two representative paradigms in misinformation detection, distinguished by their reasoning focus: (1) The implication-oriented (I) approach, which focuses on the inferred implications of the news sample; and (2) The consistency-oriented (C) approach, which assesses consistency both across modalities and between the image–caption pair and the reference article.
-Unless otherwise specified, all experiments on DeceptionDecoded default to the consistency-oriented (C) approach.
-Experimental setup and evaluation prompt details are shown in Appendix C.1.
-4.2 Main Results
-Consistency-Oriented Reasoning Generally Outperforms Implication-Oriented Reasoning.
-As shown in Table 2, a possible reason is that misleading news is often constructed through deliberate content manipulation.
-Detecting unsubstantiated inconsistencies, either between the image and its caption or between the image-caption pair and the reference article, helps models better identify misleading intent.
+Table 3: Performance (F1%) on creator desire inference. Green cells indicate the best-performing models, and red cells zero performance.
+Model M-Sub (Text) M-Sig (Text)
+I+T T+A Full I+T T+A Full
+GPT-4o 21.8 63.4 70.2 51.6 88.4 93.3
+GPT-4o-mini 3.6 36.8 18.8 10.2 64.4 45.9
+Qwen2.5-72B 10.6 53.6 49.6 24.6 83.8 82.5
+Qwen2.5-32B 6.2 35.8 27.5 18.4 67.2 58.6
+Table 4: VLM performance (Acc.%) on misleading intent prediction under partial modality settings.
+Results suggest that smaller VLMs tend to over-rely on image–text consistency when detecting misleading intent.
+(I+T: image & caption only; T+A: caption & reference article only; Full: all modalities.)
+4.3 VLMs Struggle Beyond Surface-Level Signals
+VLMs Can Be Misled by Image–Text Semantic Consistency.
+Although misleading intent is often unsupported when compared with the trustworthy news context, the manipulated image and caption typically appear internally consistent (Figure 1).
+Ideally, models should integrate information across all three modalities: image, text, and article, to identify such deception.
+Figure 3: Misleading intent detection performance of VLMs on misleading text presented either in a professional tone (as used in the original DeceptionDecoded setting) versus an explicitly misleading style.
+To investigate this, we evaluate VLMs under two partial input settings: (1) Image + Text, and (2) Text + Article.
+As shown in Table 4, the stronger GPT-4o model follows the expected trend, achieving its best performance when all modalities are available.
+In contrast, smaller VLMs deviate from this ideal.
+For example, GPT-4o-mini performs better when detecting misleading text by comparing it directly with the article, yet its performance drops notably once the misleading image is added.
+This suggests that internal image–text consistency can mask deception, exposing a weakness in models that lack deeper reasoning about the credibility of information across modalities.
+VLMs Fail to Detect Misleading Text Framed in Credible-Sounding Styles.
+Prior work on textual misinformation detection (Chen & Shu, 2024; Wu et al., 2024) shows that large language models (LLMs) often rely on stylistic cues rather than factual correctness when making veracity judgments.
+To test whether this bias extends to the multimodal setting, we reframe misleading captions in DeceptionDecoded from professional, authoritative tones to more explicitly deceptive or sensational ones (see details in Appendix F.2).
+As illustrated in Figure 3, we observe the same trend: VLMs are more likely to misclassify misleading content when it is presented in a polished, credible style.
+This suggests that current models may be disproportionately influenced by surface-level linguistic cues, posing serious risks when misinformation is deliberately crafted to appear authoritative.
 Model All Misleading (Acc. %) All Non-Misleading (Acc. %)
@@
 Qwen2.5-VL-32B 35.3 99.5 (+64.2) 1.2 (-34.1) 95.9 99.6 (+3.7) 3.4 (-92.5)
-Table 4: Misleading intent detection performance of VLMs when exposed to spurious authenticity-related hints, averaged over both misleading and non-misleading cases.
-Hint construction details are provided in Section 5.1.
-Model M-Sub (Text) M-Sig (Text)
-I+T T+A Full I+T T+A Full
-GPT-4o 21.8 63.4 70.2 51.6 88.4 93.3
-GPT-4o-mini 3.6 36.8 18.8 10.2 64.4 45.9
-Qwen2.5-72B 10.6 53.6 49.6 24.6 83.8 82.5
-Qwen2.5-32B 6.2 35.8 27.5 18.4 67.2 58.6
-Table 5: VLM performance (Acc.
-%) on misleading intent prediction under partial modality settings, suggesting that smaller VLMs may overly rely on image-text consistency for misleading intent detection.
-(I+T: image & caption only; T+A: text & reference article only; Full: all modalities provided.)
-Figure 3: Misleading intent detection performance of VLMs on misleading text presented either in a professional tone (as used in the original DeceptionDecoded setting) or in an explicitly misleading style.
-VLMs Struggle to Reason About Misleading Creator Intent.
-As shown in Table 2, even state-of-the-art MLRMs such as Claude-3.7 demonstrate limited performance in detecting and attributing misleading creator intent.
-Performance is even lower on the more complex task of creator desire inference (Table 3), suggesting a deeper challenge in reasoning about communicative goals.
-These findings raise alarming concerns.
-Our setup reflects a realistic deployment scenario, where readers are exposed to subtle yet misleading representations of otherwise trustworthy multimodal news and must rely on automated systems for guidance.
-The consistent underperformance of VLMs under such conditions suggests an urgent need to investigate the underlying patterns these models exploit when assessing content authenticity, inspiring our in-depth analysis in Section 5.
-5 Analysis
-5.1 Probing the Limitations of VLMs
-VLMs Can Be Misled by Surface-Level Image–Text Consistency.
-Although misleading intent is often not substantiated when compared with the trustworthy reference article, the manipulated image and caption typically appear internally consistent, as illustrated in Figure 1.
-Ideally, models should integrate information across all modalities (i.e., image, text, and article) to detect such deceptive intent.
-To investigate this, we evaluate VLMs under two partial input settings: (1) Image + Text, and (2) Text + Article.
-However, results in Table 5 reveal a contradiction to this ideal.
-For instance, GPT-4o-mini performs better when detecting misleading text by comparing it with the article, but its performance drops notably when the misleading image is added.
-This suggests that consistency between image and text can mask deception, misleading models that lack deeper reasoning about cross-source veracity.
-VLMs Fail to Detect Misleading Text Framed in Credible-Sounding Styles.
-Prior work in textual misinformation detection Chen and Shu (2024); Wu et al.
-(2024) has shown that large language models (LLMs) often rely on stylistic cues rather than factual correctness when making veracity judgments.
-To investigate whether this stylistic bias extends to the multimodal setting, we reframe misleading captions in DeceptionDecoded from professional-sounding styles to more overtly deceptive or sensational tones (see Appendix C.2).
-As illustrated in Figure 3, we observe a similar trend: VLMs are more likely to misclassify misleading content when it is framed in a polished, authoritative style.
-This suggests that current models may be overly influenced by surface-level linguistic cues, posing a serious risk when misinformation is intentionally crafted to appear credible.
-Models React Strongly to Spurious Authenticity Cues in Prompts.
-To examine whether VLMs are influenced by high-level framing cues that are not grounded in the input content, we inject two types of authenticity-related hints into the prompt (see Appendix C.3): (1) Skeptical, which presumes the news piece may contain intentional distortions.
-This hint serves as a helpful signal for misleading samples but becomes adversarial for non-misleading samples; and (2) Trusting, which assumes the content is reliable, reversing the helpful and adversarial roles.
-As shown in Table 4, these prompts produce significant and consistent performance changes: they improve accuracy in helpful cases and degrade performance in adversarial ones.
-This reveals that VLMs are susceptible to spurious prior assumptions about content authenticity, even when such cues are not supported by the input.
-5.2 Impact of Image Generation Advances
-DeceptionDecoded includes reasonably high-quality images generated using the FLUX model, as supported by human validation results in Section 3.4.
-As image generation capabilities continue to advance, we conduct a pilot study to assess whether VLM performance improves when exposed to higher-fidelity images generated by GPT-image-1 OpenAI (2025a), using a subset of 200 sampled misleading instances.
-As shown in Table 6 and Figure 4, we observe modest performance gains, likely due to the more vivid visual portrayal of creator intent.
-However, despite these improvements, understanding the underlying communicative intent behind multimodal news remains a significant challenge even when image realism is enhanced.
-These findings highlight that surface-level fidelity alone is insufficient, reinforcing the need for models that reason explicitly about creator objectives.
-The growing realism of deceptive images as reflected by GPT-generated instances introduces a significant risk.
-When high-fidelity images convincingly align with misleading textual narratives, the deception becomes more subtle and harder to detect.
-These advances further underscore the importance of reasoning about the creator’s underlying intent, rather than relying solely on content realism or modality alignment.
-Figure 4: Case Study: a VLM (GPT-4o-mini) fails to detect misleading creator intent in a FLUX-generated image due to its inability to recognize the unsubstantiated presence of a crowd with fireworks, given the news context.
-The state-of-the-art GPT-image-1 model produces a clearer depiction, leading to a correct prediction.
+Table 5: VLM performance on misleading intent detection with spurious authenticity hints, averaged over misleading and non-misleading cases.
+Helpful / Adversarial hint construction details are provided in Section 4.3.
+VLMs React Strongly to Spurious Authenticity Cues in Prompts.
+We further test whether models are influenced by high-level framing cues that are not grounded in the input content.
+To do so, we inject two types of authenticity-related hints into the prompts (see Appendix F.3): (1) Skeptical, which presumes the news piece may contain intentional distortions (helpful for misleading samples but adversarial for non-misleading ones), and (2) Trusting, which assumes the content is reliable (reversing the helpful/adversarial roles).
+As shown in Table 5, these hints produce significant and systematic performance shifts: they improve accuracy in helpful cases but sharply degrade it in adversarial ones.
+This demonstrates that VLMs remain highly sensitive to spurious prior assumptions about authenticity, even when such cues are not supported by the input.
+5 Discussion
+5.1 Broader Utility of Creator Intent Identification
+Model Macro-F1 (%) Macro-Acc. (%)
+Llava-v1.6-7B 44.41 50.31
+   + FFT (w/ 6k samples) 52.37 (+7.96) 63.52 (+13.21)
+Qwen2.5-VL-7B 27.96 52.12
+   + FFT (w/ 6k samples) 58.66 (+30.70) 67.24 (+15.12)
+Table 6: Further utility of DeceptionDecoded in enhancing multimodal misinformation detection (MMD).
+Fine-tuning VLMs on DeceptionDecoded yields consistent performance improvements on MMFakeBench, a general MMD benchmark.
+A central contribution of DeceptionDecoded is its focus on misleading creator intent, an aspect largely overlooked in existing multimodal misinformation benchmarks yet essential for developing robust MMD systems.
+To examine its broader utility, we conduct a transfer experiment on the MMFakeBench validation set (Liu et al., 2025), which contains 1,000 samples (700 fake, 300 real).
+Specifically, we compare the performance of LLaVA-v1.6-7B and Qwen-2.5-VL-7B before and after full-parameter fine-tuning on 6,000 DeceptionDecoded samples (1,000 per class across six intent categories, with misleading samples grouped as “Fake” and non-misleading as “Real”) for one epoch.
+The setup is detailed in Appendix F.5.
+As shown in Table 6, fine-tuning yields consistent and significant performance gains across metrics, indicating that DeceptionDecoded not only facilitates intent-based modeling but also enhances general MMD.
+5.2 Image Realism Outpaces VLMs’ MMD Capabilities
+DeceptionDecoded incorporates reasonably high-quality images generated with FLUX.1 [dev] (Black Forest Labs, 2024), as confirmed by human validation in Section 3.4 and Appendix E.
+To examine the implications of recent advances in image generation, we conducted a controlled evaluation using a subset of 200 sampled misleading instances, replacing FLUX outputs with higher-fidelity versions generated by GPT-image-1 (OpenAI, 2025b).
 Model M-Sub (Image) M-Sig (Image)
@@
 Qwen2.5-32B 30.0 37.0 46.5 46.0
-Table 6: Misleading intent detection performance of VLMs on images created using different image generation models, evaluated on a pilot test set of 200 samples.
-While performance improves with higher-fidelity GPT images, a substantial gap remains.
+Table 7: Misleading intent detection performance of VLMs on images generated by different models. Evaluation is conducted on a 200-sample pilot set.
+As shown in Table 7 and Figure 4, several VLMs exhibit modest, consistent gains when evaluated on GPT-generated images, likely because higher visual fidelity clarifies intent-related cues (for example, more legible, vivid depictions of fireworks and crowds than the FLUX-generated variant).
+These observations indicate that stronger visuals can make certain manipulations easier to detect.
+However, despite these improvements, modeling the underlying communicative intent in multimodal news remains difficult.
+Surface-level fidelity is insufficient; effective detection requires reasoning explicitly about creator objectives.
+At the same time, the growing realism of GPT-generated images introduces substantial risks.
+When high-fidelity visuals align convincingly with misleading textual narratives, the resulting deception becomes more subtle and more difficult to detect.
+In our experiments, prompting GPT-image-1 to simulate malicious creators, without explicitly requesting misinformation, is able to bypass the established safety guardrail and generate convincing manipulations for all 200 cases.
+The combination of scalability and realism heightens risk, particularly since VLMs remain weak at intent detection even under these enhanced conditions.
+Figure 4: Case Study: a VLM (GPT-4o-mini) fails to detect misleading creator intent in a FLUX-generated image.
+The model overlooks the presence of a crowd with fireworks, an unsubstantiated addition given the powerhouse fire described in both the news context and the caption.
+In contrast, the state-of-the-art GPT-image-1 model produces a more vivid depiction of fireworks, enabling the correct prediction.
 6 Conclusion
-In this work, we establish the important yet underexplored task of understanding creator intent in multimodal misinformation.
-We introduce DeceptionDecoded, a large-scale benchmark designed to evaluate VLMs across intent-centric tasks.
-Through extensive experiments, we uncover alarming vulnerabilities in state-of-the-art VLMs, which over-rely on spurious shortcuts such as cross-modal consistency, stylistic cues, and heuristic indicators of authenticity to detect misleading intent.
-Our findings pave the way for future research on MMD systems capable of intent-aware, authenticity-centric semantic reasoning toward effective misinformation governance.
-Limitations
-Through DeceptionDecoded, we establish a framework for understanding misleading creator intent in multimodal news, providing a meaningful first step toward modeling and detecting misleading creator intent.
-While DeceptionDecoded presents a structured and scalable framework, it also opens up several avenues for further research in advancing intent-aware misinformation governance across more realistic and socially grounded scenarios.
-First, although our framework supports the evaluation of creator intent interpretation through tasks such as misleading intent detection and creator desire inference, fully open-ended intent articulation and corresponding evaluation remains an open challenge.
-Future work could explore adaptive prompting strategies and leverage the emerging LLM-as-a-Judge paradigm Gu et al.
-(2024) to assess generated intent representations in more flexible and scalable ways.
-Second, our current benchmark focuses on intent detection and interpretation, but does not directly measure downstream societal influence.
-In real-world scenarios, the effectiveness of misinformation depends not only on creator intent but also on how audiences interpret and react to the content.
-Incorporating social grounding, such as user feedback signals, comment chains, or engagement patterns on social media, would provide richer context for evaluating the influence of misleading content and calibrating model responses accordingly.
-Despite these limitations, DeceptionDecoded represents an important contribution toward systematic, intent-aware evaluation of VLMs in real-world misinformation contexts.
-We hope our benchmark and insights will serve as a solid foundation for future work.
-Ethics Statement
-The intent-guided simulation of multimodal misinformation in DeceptionDecoded involves strategies that could potentially be misused to generate misleading content using VLMs and open-source image generation models (e.g., FLUX).
-Nonetheless, given the realistic and challenging nature of detecting misleading creator intent in multimodal news, we believe it is important to transparently acknowledge these risks while reporting empirical findings on the performance and limitations of current state-of-the-art VLMs.
-Our benchmark is developed solely to highlight the important yet underexplored aspect of modeling malicious creator intent in MMD, and to advance the understanding of current system limitations.To minimize harm, we explicitly avoid manipulating news content involving specific, identifiable individuals; all simulated content is based on anonymized or synthetic entities.
-To further mitigate misuse, we will open-source the dataset and evaluation scripts, but refrain from releasing the specific generation prompts that could be repurposed for deceptive use.
-Data access will be granted only to verified researchers under a binding usage agreement to ensure responsible use.
-All data collection and usage comply with the terms of service of the underlying models and platforms.
-Appendix A DeceptionDecoded Details
-A.1 News Topics
-VisualNews Liu et al.
-(2021) provides fine-grained topic annotations for each news article.
+We establish creator intent as a central perspective for multimodal misinformation and introduce DeceptionDecoded, a large-scale benchmark grounded in trustworthy news contexts.
+Through extensive experiments on 14 VLMs, we reveal that current state-of-the-art models are weak at intent reasoning, often relying on spurious cues such as cross-modal consistency, stylistic polish, and heuristic authenticity signals.
+These findings demonstrate that advances in visual fidelity are insufficient on their own, highlighting the need for authenticity-centric, intent-aware modeling.
+Taken together, DeceptionDecoded offers not only an evaluation benchmark but also frames a research agenda, opening new directions for intent-aware evaluation and governance of multimodal misinformation at scale.
+[^1] Limitations and ethical considerations are discussed in Appendix A and Appendix B.
+Appendix A Limitations and Future Work
+Through DeceptionDecoded, we establish a structured and scalable framework for understanding misleading creator intent in multimodal news, representing a meaningful first step toward intent-aware multimodal misinformation governance.
+While our benchmark makes a substantial contribution, it also opens several promising directions for extending its value and impact.
+First, although DeceptionDecoded supports intent interpretation through tasks such as misleading intent detection and creator desire inference, fully open-ended intent articulation and evaluation remains an open challenge.
+Future work could explore adaptive prompting strategies or leverage emerging paradigms such as LLM-as-a-Judge (Gu et al., 2024) to assess generated intent representations in more flexible, reliable, and scalable ways.
+Second, while our current focus is on intent detection and interpretation, the downstream societal influence of misinformation is not directly measured.
+In real-world settings, the effectiveness of misleading narratives depends not only on creator intent but also on how audiences interpret and respond.
+Incorporating social grounding (such as user feedback signals, comment chains, or engagement patterns) would provide richer context for evaluating the reach and impact of misinformation, and for calibrating model responses accordingly.
+Overall, DeceptionDecoded provides a principled foundation for systematic, intent-aware evaluation of VLMs, while also charting a path toward more socially grounded, resilient, and scalable approaches for governing multimodal misinformation.
+Appendix B Ethics Statement
+The intent-guided simulation of multimodal misinformation in DeceptionDecoded necessarily involves strategies that could be misused to generate misleading content with VLMs and open-source image generation models (e.g., FLUX).
+Nonetheless, given the pressing challenge of detecting misleading creator intent in multimodal news, it is important to transparently acknowledge these risks while reporting empirical findings on the strengths and limitations of current state-of-the-art VLMs.
+Our benchmark is developed solely to highlight the important yet underexplored aspect of modeling malicious creator intent in MMD, and to advance the understanding of current system limitations.
+To minimize harm, we explicitly avoid manipulating news content involving specific, identifiable individuals; all simulated content is based on anonymized or synthetic entities.
+To further mitigate misuse, we will open-source the dataset and evaluation scripts but refrain from releasing the specific generation prompts that could be repurposed for deception.
+Access will be restricted to verified researchers under a binding usage agreement, and all data collection complies with the terms of service of the underlying models and platforms.
+Overall, we prioritize transparency while enforcing safeguards against misuse.
+Importantly, our work does not seek to promote or enforce a single interpretation of news events.
+Instead, it focuses on detecting creator intent when multimodal content distorts the implications of a trustworthy news context by embedding unsubstantiated claims or implications (see example in Figure 1 and Figure 2).
+This framing distinguishes between responsible reporting, which preserves neutrality, and misleading reporting, which introduces unsupported interpretations.
+Ultimately, our goal is to support the development of intent-aware systems that safeguard trustworthy reporting while respecting the rights of information consumers to hold diverse viewpoints.
+Appendix C LLM Usage Statement
+LLMs were used solely for minor grammatical editing of the manuscript. All methodological design, analysis, and writing were conducted under full human oversight and responsibility.
+Appendix D DeceptionDecoded Details
+D.1 News Topics
+VisualNews (Liu et al., 2021) provides fine-grained topic annotations for each news article.
 For DeceptionDecoded, we select 10 major topic categories based on sample abundance, topical diversity, and real-world relevance.
@@
 These topics form the foundation for generating diverse and realistic multimodal misinformation instances in DeceptionDecoded.
-A.2 Source Data Acquisition
+D.2 Source Data Acquisition
 In Section 3.1, we obtain high-quality source data for DeceptionDecoded with the prompt outlined in Figure 5.
-A.3 Aspects of Desired Influence
-We define a set of major desired aspects of societal influence to facilitate creator intent establishment in Section 3.2. The aspects are listed as follows:
-Political Polarization, Social Polarization, Cultural and Religious Polarization, Economic Misinformation, Public Health and Safety, Environmental and Scientific Polarization, Geopolitical and International Relations, Psychological and Emotional Manipulation
-A.4 DeceptionDecoded Examples
-Figure 5: An example in DeceptionDecoded, showcasing a scenario where the creator intentionally misrepresents the news content with a distorted caption to mislead readers.
-Figure 6: An example in DeceptionDecoded, showcasing a scenario where the creator intentionally misrepresents the news content with a distorted image to mislead readers.
-We present two DeceptionDecoded data examples in Figure 5 (textual modification) and Figure 6 (visual modification).
-To prevent misuse, we refrain from disclosing the data generation prompts for Section 3.3 that could be repurposed for deceptive use.
-Appendix B Human Evaluation Details
-We illustrate the annotation process for human evaluation in Appendix B in Figure 7.
-Figure 7: Illustration of human evaluation instructions for validating the quality of DeceptionDecoded.
-Appendix C VLM Evaluation Details
-C.1 Experimental Setup
-With DeceptionDecoded, we systematically evaluate 14 representative VLMs. Detailed model names and versions are provided in Table 7.
-For all evaluations, we set the model temperature to 0 to ensure stable outputs, and run each experiment once.
-Detailed evaluation prompts for the implication- and consistency-based approaches are shown in Appendix C.4.
-C.2 Caption Style Reframing Formulation
-We adopt the following prompt for the stylistic reframing experiments in Section 5.1 and Table 3.
-You are given a factual news caption.
-Your task is to rewrite it in an unreliable style: that is, make it sound sensational, biased, emotional, or manipulative, while preserving the original content and factual meaning.
-Do not add or remove any factual information.
-Only change the tone and style to reflect how misinformation might manipulate readers through language.
-Original Caption: {caption} Unreliable-style Caption:
-C.3 Authenticity-Related Hint Formulation
-To investigate how VLMs respond to subtle authenticity cues, we design two types of prompt-level hints aligned with the analysis in Section 5.1.
-These hints are inserted into the input to simulate prior assumptions about content reliability.
-Hint (Skeptical): “Hint: The news piece is likely to contain intentional distortions. Carefully analyze the image and caption to identify any misleading or manipulative elements.”
-Hint (Trusting): “Hint: The news piece is sourced from reputable outlets and is therefore unlikely to contain intentional distortions. ”
-C.4 Evaluation Prompts
-We present our prompts for evaluating VLMs in Figure 9 (implication-oriented) and Figure 10 (consistency-oriented).
 Data filtering in DeceptionDecoded You are a senior journalist tasked with evaluating a news caption based on five key criteria: - Public Interest: Does the caption describe a significant event likely to capture public attention?
@@
 Caption: {caption} Response:
-Figure 8: Prompt for selecting VisualNews Liu et al. (2021) samples as source data for DeceptionDecoded.
+Figure 5: Prompt for selecting VisualNews (Liu et al., 2021) samples as source data for DeceptionDecoded.
+D.3 Aspects of Desired Influence
+We define a set of major desired aspects of societal influence to facilitate creator intent establishment in Section 3.2. The aspects are listed as follows:
+Political Polarization, Social Polarization, Cultural and Religious Polarization, Economic Misinformation, Public Health and Safety, Environmental and Scientific Polarization, Geopolitical and International Relations, Psychological and Emotional Manipulation
+D.4 DeceptionDecoded Data Examples
+To illustrate how intent-guided modifications manifest in practice, we present two representative DeceptionDecoded examples: a textual modification in Figure 6, where the caption is manipulated to convey misleading intent while the image remains unchanged, and a visual modification in Figure 7, where the image is altered to mislead while the caption is preserved.
+These examples highlight how creator intent can be embedded through either modality, reinforcing the multimodal nature of the benchmark.
+Appendix E Human Evaluation Details
+DeceptionDecoded benefits from our proposed synthetic framework for intent-guided multimodal news creation, which grounds manipulations in trustworthy news contexts while explicitly modeling creator intent.
+To validate that the generated data are realistic, relevant, and aligned with intended objectives, we conduct a series of human evaluation studies.
+These evaluations provide assurance that the benchmark is both reliable for systematic model testing and meaningful for real-world misinformation scenarios.
+Beyond the central evaluation of misleading vs.
+non-misleading perception in Section 3.4 (detailed in Appendix E.1), we further assess two additional aspects: data realism (Appendix E.2) and intent alignment (Appendix E.3).
+Across all studies, three independent graduate student annotators followed the same setup described in Section 3.4.
+Accuracy is reported based on majority-vote labels, and inter-annotator agreement is measured using Fleiss’ $\kappa$.
+Together, these evaluations provide a comprehensive view of data validity from multiple perspectives.
+E.1 Human Evaluation of Data Quality
+This evaluation (results detailed in Section E) establishes whether human annotators perceive the generated manipulations as misleading or non-misleading in line with the intended labels.
+Validating this perception is essential because it ensures that the benchmark captures manipulative strategies in ways that are interpretable to human readers, not just detectable by models.
+We illustrate the annotation process in Figure 8.
+Figure 6: An example in DeceptionDecoded, showcasing a scenario where the creator intentionally misrepresents the news content with a distorted caption to mislead readers.
+Figure 7: An example in DeceptionDecoded, showcasing a scenario where the creator intentionally misrepresents the news content with a distorted image to mislead readers.
+Figure 8: Human evaluation instructions for assessing the data quality in DeceptionDecoded (detailed in Section 3.4).
+E.2 Human Evaluation of Data Realism
+While label accuracy validates whether manipulations are recognized as misleading, it is equally important to test whether the content feels plausible to humans.
+If generated samples are unrealistic, then models trained on them would not generalize to real-world misinformation scenarios.
+Standalone Plausibility.
+Setup: Annotators judged whether a synthesized news piece could plausibly appear in today’s news or on social media.
+Question: “After reading the synthesized news piece, could this plausibly appear on today’s news or social media?”
+  • 120 text-synthesized samples (80 misleading, 40 non-misleading): 98.3% accuracy, $\kappa=0.854$.
+  • 120 image-synthesized samples (80 misleading, 40 non-misleading): 91.7% accuracy, $\kappa=0.815$.
+High standalone plausibility confirms that samples resemble authentic news-style reporting and could realistically challenge automated systems.
+Pairwise Plausibility.
+Setup: Annotators compared original $\rightarrow$ manipulated pairs and judged whether the manipulated version could plausibly appear as deceptive reporting.
+Question: “Given the original $\rightarrow$ manipulated pair, could the manipulated version plausibly appear on today’s news or social media, created by an untrustworthy publisher to mislead the audience?”
+  • 80 misleading text-synthesized samples: 97.5% accuracy, $\kappa=0.853$.
+  • 80 misleading image-synthesized samples: 87.5% accuracy, $\kappa=0.663$.
+Strong pairwise plausibility shows that manipulations are subtle enough to deceive when contrasted with authentic reporting, reinforcing their utility for evaluating fine-grained detection capabilities.
+E.3 Human Evaluation of Intent Alignment
+Finally, beyond realism and label validity, we test whether manipulations remain faithful to the intended creator objectives (Section 3.2).
+Verifying intent alignment is necessary, since without alignment to the desired influence or execution plan, the dataset would fail to capture the communicative strategies central to misinformation governance.
+Desire Alignment.
+Setup: Annotators judged whether manipulations matched the creator’s specified societal target (e.g., political polarization, public health).
+Question: “After reviewing the original $\rightarrow$ manipulated pair, is the manipulation consistent with the creator’s established desire to exert negative influence on the specified societal aspect(s): {list of desire aspects}?”
+  • 80 misleading text-synthesized samples: 92.5% accuracy, $\kappa=0.732$.
+  • 80 misleading image-synthesized samples: 86.3% accuracy, $\kappa=0.636$.
+Execution Plan Alignment.
+Setup: Annotators judged whether manipulations followed the specified execution plan of the creator.
+Question: “After reviewing the original $\rightarrow$ manipulated pair, is the manipulation consistent with the creator’s established execution plan: {plan}?”
+  • 80 misleading text-synthesized samples: 93.8% accuracy, $\kappa=0.772$.
+  • 80 misleading image-synthesized samples: 86.3% accuracy, $\kappa=0.643$.
+These evaluations confirm that manipulations are not only plausible but also systematically aligned with simulated intent.
+This alignment ensures that DeceptionDecoded captures the communicative mechanisms underlying misinformation, making it a robust benchmark for intent-aware multimodal misinformation detection.
+Appendix F VLM Evaluation Details
+Model Model Card
+o4-mini (OpenAI, 2025a) o4-mini-2025-04-16
+Claude-3.7-Sonnet (Anthropic, 2025b) claude-3-7-sonnet-20250219
+Gemini-2.5-Pro (Google, 2025) gemini-2.5-pro-preview-03-25
+GPT-4o (Hurst et al., 2024) gpt-4o-2024-08-06
+Claude-3.5-Sonnet (Anthropic, 2025a) claude-3-5-sonnet-20241022
+Gemini-1.5-Pro (Georgiev et al., 2024) gemini-pro-1.5
+GPT-4o-mini (OpenAI, 2024) gpt-4o-mini-2024-07-18
+Claude-3.5-Haiku (Anthropic, 2025c) claude-3-5-haiku-20241022
+Qwen2.5-VL-72B-Instruct (Qwen Team, 2025) qwen2.5-vl-72b-instruct
+Qwen2.5-VL-32B-Instruct (Qwen Team, 2025) qwen2.5-vl-32b-instruct
+LLama-3.2-Vision-Instruct-11B (Grattafiori et al., 2024) llama-3.2-11b-vision-instruct
+InternVL2-8B (Chen et al., 2024) InternVL2-8B
+LLaVA-1.6-7B (Liu et al., 2024) llava-v1.6-mistral-7b
+Qwen2.5-VL-7B-Instruct (Qwen Team, 2025) Qwen2.5-VL-7B-Instruct
+Table 8: Model cards for all VLMs evaluated throughout our work.
+F.1 Experimental Setup
+With DeceptionDecoded, we systematically evaluate 14 representative VLMs spanning different families, sizes, and access levels.
+Detailed model names and versions are listed in Table 8.
+To ensure reproducibility and stability, all evaluations are run once with temperature set to 0.
+This removes sampling variance and allows fair comparisons across models.
+Evaluation prompts for the implication-oriented and consistency-oriented approaches are provided in Appendix F.4.
+F.2 Caption Style Reframing Formulation
+Stylistic cues can strongly influence model predictions even when the underlying factual content is unchanged.
+To probe this vulnerability, we reframe original captions in DeceptionDecoded—which adopt a professional, neutral news tone—into manipulative styles that sound sensational, biased, or emotional.
+This allows us to test whether models mistake stylistic credibility for factual reliability.
+The experiments in Section 4.3 and Table 3 examine whether VLMs over-rely on linguistic polish or credibility markers instead of reasoning about creator intent and cross-modal consistency at implication level.
+The exact prompt used for stylistic reframing is shown below:
+You are given a factual news caption.
+Your task is to rewrite it in an unreliable style: that is, make it sound sensational, biased, emotional, or manipulative, while preserving the original content and factual meaning.
+Do not add or remove any factual information.
+Only change the tone and style to reflect how misinformation might manipulate readers through language.
+Original Caption: {caption} Unreliable-style Caption:
+Table 9 shows side-by-side comparisons of professional (default in DeceptionDecoded) vs. unreliable styles, highlighting how superficial stylistic changes can shift perceived reliability.
+Professional Style Unreliable Style
+A firefighter struggles alone against a massive wildfire due to severe budget cuts, highlighting the government’s failure to support emergency services.
+A lone firefighter heroically battles a monstrous wildfire, abandoned by a negligent government that has slashed emergency service budgets, leaving our communities to burn!
+Massive explosion at a secret military facility near Santa Clarita raises questions about government transparency.
+Shocking blast at hidden military base near Santa Clarita sparks outrage and demands for government accountability!
+Terrorist attack causes massive fire on M6, authorities urge residents to stay indoors.
+Catastrophic terrorist attack ignites inferno on M6, authorities desperately plead with terrified residents to hide indoors!
+Table 9: Examples contrasting professional and unreliable reporting styles.
+F.3 Authenticity-Related Hint Formulation
+Prompt-level framing can introduce subtle biases that shape model predictions even before any multimodal evidence is processed.
+To probe this vulnerability, we inject authenticity-related hints that explicitly bias models toward or against assuming reliability.
+We design two types of hints, described in Section 4.3, and insert them directly into the evaluation prompts to simulate prior assumptions about content trustworthiness.
+This setup allows us to test whether VLMs rely on ungrounded framing cues rather than reasoning from the actual news content.
+The two hint variants are defined as follows:
+Hint (Skeptical): “Hint: The news piece is likely to contain intentional distortions. Carefully analyze the image and caption to identify any misleading or manipulative elements.”
+Hint (Trusting): “Hint: The news piece is sourced from reputable outlets and is therefore unlikely to contain intentional distortions. ”
+F.4 Evaluation Prompts
+In Section 4.2, we evaluate 14 VLMs on the core task of detecting misleading creator intent.
+The full evaluation prompts are illustrated in Figure 9 (implication-oriented reasoning) and Figure 10 (consistency-oriented reasoning).
 Implication-Oriented Approach (I) Your task is to determine whether a news piece has been intentionally distorted to mislead readers.
@@
 Do not add any extra line breaks or formatting outside this structure: { “pred”: “Answer to Q1, followed by a brief explanation”, “source”: "Answer to Q2", “desire”: "Answer to Q3" }
-Figure 9: Prompt for the implication-oriented approach (I) for detecting misleading news creator intent.
+Figure 9: Prompt for the implication-oriented (I) approach to detecting misleading creator intent, utilized in Section 4.2.
 Consistency-Oriented Approach (C) Your task is to determine whether a news piece has been intentionally distorted to mislead readers.
@@
 Do not add any extra line breaks or formatting outside this structure: { “pred”: “Answer to Q1, followed by a brief explanation”, “source”: "Answer to Q2", “desire”: "Answer to Q3" }
-Figure 10: Prompt for the consistency-oriented approach (C) for detecting misleading news creator intent.
+Figure 10: Prompt for the consistency-oriented (C) approach to detecting misleading creator intent, utilized in Section 4.2.
+F.5 VLM Fine-Tuning Setup
+As described in Section 5.1, we evaluate the broader utility of creator intent identification for general multimodal misinformation detection by fully fine-tuning two open-source VLMs on DeceptionDecoded: LLaVA-v1.6-Mistral-7B and Qwen-2.5-VL-7B.
+Each model is trained for one epoch on 6,000 DeceptionDecoded samples (six intent classes, 1,000 per class) and then tested on MMFakeBench (Liu et al., 2025), a general MMD benchmark.
+For evaluation, we use MMFakeBench’s validation split of 1,000 instances (700 fake, 300 real).
+To prevent label-order bias, the training samples are randomly shuffled, since the raw files list all non-misleading cases first and all misleading cases last.
+Each training instance consists of an image paired with its caption, inserted into a fixed prompt (shown verbatim in Figure 11, with {caption} substituted).
+The model produces a binary “Yes/No” prediction, where “Yes” indicates misleading and “No” indicates non-misleading.
+Supervision is answer-only: all prompt tokens are masked, and the loss is computed solely on the terminal answer span, located using a robust matcher tolerant to whitespace and punctuation variations.
+Inputs are left-padded and encoded without truncation to preserve full context.
+Optimization employs an effective batch size of 32 (per-device batch size 1 with gradient accumulation 32), a learning rate of $1\times 10^{-5}$, and bf16 mixed precision.
+This configuration is designed to isolate the decision boundary, align the training objective with the evaluation setting, and maintain consistent training capacity across both model families.
+Prompt for VLM Fine-Tuning You are given a piece of multimodal news content consisting of an image and its caption: {caption}.
+Determine whether this image–caption pair contains misinformation.
+Specifically: 1.
+Identify any factual inaccuracies in the caption.
+2.
+Check for inconsistencies or mismatches between the image and the caption.
+Answer “Yes” if the content contains misinformation, or “No” if it does not.
+Important: Focus only on the news content itself.
+The presence of AI-generated artifacts or stylistic cues alone does not indicate misinformation.
+The answer is:
+Figure 11: Prompt for the VLM fine-tuning experiments in Section 5.1 and Appendix F.5.